### v1

# hep-th/yymmnnn

Abstract
In this paper, we study black hole solutions in double field theory. In the first part, we introduce a solution-generating method and classify black hole solutions into three categories in standard double field theory. To solve the problem of double time, we utilize space/time split double field theory in the second part to derive black hole solutions. By introducing a cosmological constant and imposing the strong constraint, our findings indicate that when considering the entire doubled spacetime, the curvature signifies an AdS vacuum. However, in the subspace of the entire spacetime, black holes and curvature singularities emerge.
1 Introduction
Double field theory (DFT) is a recent progress in closed string theory. This approach doubles all spacetime coordinates and elevates T-duality to an intrinsic symmetry of the action [1, 2, 3, 4, 5]. In closed string theory, when a string moves in compactified spacetime, the momentum of its center of mass can be quantized and is conjugate to the coordinates $x^{i}$. Additionally, if the string wraps around the compactified background multiple times, it introduces another quantum number, the winding number. Consequently, the physical observables of the string depend on both the momentum and the winding number. T-duality, which exchanges momentum and winding numbers between two strings wrapped on different compactified radii, requires treating these two quantum numbers equally within a field theory. Thus, it is reasonable to introduce extra coordinates $\tilde{x}_{i}$ conjugate to the winding number. DFT suggests that the usual spacetime coordinates $x^{i}\left(i=1,2,\cdots,D\right)$ can be doubled to $X^{M}=\left(\tilde{x}_{\mu},x^{\mu}\right)$, where $M=1,2,\cdots,2D$. Consequently, all spacetime component fields, including the gravitational field $g_{\mu\nu}$, the dilaton $\phi$, and the anti-symmetric Kalb-Ramond field $b_{\mu\nu}$, depend on $X^{M}$. In this framework, $g_{\mu\nu}$ and $b_{\mu\nu}$ combine into a generalized metric that transforms linearly under continuous $O\left(D,D,\mathbb{R}\right)$ transformations. The generalized coordinate transformations are also well-defined on this $2D$ manifold. Based on these features, the DFT action can be constructed to explicitly exhibit $O\left(D,D,\mathbb{R}\right)$ symmetry. Moreover, DFT must satisfy the strong constraint, meaning the theory lives in a $D$-dimensional subspace of the $2D$ manifold. This constraint, derived from the level-matching condition of closed string theory, requires that all fields depend on only one set of the double coordinates $\left(\tilde{x},x\right)$.
In previous works [6, 7, 8, 9, 10, 11, 12], researchers have explored the cosmological solutions of DFT. These studies have provided a consistent method to group T-dual cosmological solutions from standard string cosmology [13, 14, 15]. These solutions unify pre- and post-big bang evolutions in a single line element. Additionally, recent investigations have examined specific black hole solutions under certain constraints [16, 17, 18, 19]. However, two key issues remain unresolved:
  1. 1.
    The presence of two time coordinates in the generalized metric lacks a reasonable physical explanation. Directly imposing the strong constraint eliminates the double coordinates of the solutions, reducing them to the standard string theory solutions.
  2. 2.
    The solutions appear to unify previous T-dual solutions without introducing new physics.
Solving these two problems could make studying black hole solutions in DFT highly valuable. To tackle the first issue, we can utilize recent advances in the canonical formulation of DFT [20, 21, 22]. In this formulation, the ADM (Arnowitt-Deser-Misner)-like decomposition splits the $2D$ manifold into temporal and spatial parts $X^{M}=\left(\tilde{t},t,X^{\hat{M}}\right)$, requiring all spacetime component fields to be independent of the dual time coordinate $\tilde{t}$. Consequently, only the time coordinate remains undoubled, reducing the $2D$-dimensional DFT to $12D-$ 1 dimensions. We refer to this approach as space/time split DFT for the remainder of this paper. To illustrate how this decomposition can generate new solutions, we must clarify the differences among three types of $O\left(D,D,\mathbb{R}\right)$ invariant theories as follows:
  • Low-energy effective action with $O\left(m,m,\mathbb{R}\right)$ duality: In this action, spacetime coordinates are not doubled. The $O\left(m,m,\mathbb{R}\right)$ dduality imposes significant constraints on the form of the spacetime metric. Specifically, when the background metric is independent of $m$ coordinates, the action exhibits $O\left(m,m,\mathbb{R}\right)$ symmetry. For example, FLRW backgrounds display $O\left(D-1,D-1,\mathbb{R}\right)$ symmetry, also known as "scale-factor duality," where the action remains invariant under the scale factor transformation $a\left(t\right)\rightarrow\frac{1}{a\left(t\right)}$. Recent studies have explored the classification of all orders of $\alpha^{\prime}$ corrections based on this duality [23, 24, 25]. Non-perturbative and non-singular solutions can be found in references[26, 27, 28, 29, 30, 31, 32, 33]. Solutions involving matter sources are discussed in references [34, 35, 36, 37].
  • DFT with $O\left(D,D,\mathbb{R}\right)$ symmetry: In this formulation, all spacetime coordinates are doubled, and massless closed string fields depend on these double coordinates without needing to be independent of $m$ coordinates. The level-matching condition of closed strings imposes the strong constraint, requiring the theory to live on a $D-1$ subspace of the doubled spacetime. Vacuum cosmological solutions of DFT have been provided in [6, 7, 8] and later extended to include matter sources [9, 10, 11, 12]. Additionally, interesting solutions such as pp-waves [38], monopoles [39], and a summary of these solutions can be found in [40].
  • Space/time split DFT with $O\left(D-1,D-1,\mathbb{R}\right)$ symmetry: In this version, only the spatial coordinates are doubled, and fields are not required to be independent of $m$ coordinates. Since a metric component in the time direction is absent, solutions benefit from the double coordinates while also extending beyond simple combinations. This approach holds potential for discovering new solutions. There have been some efforts to study black hole thermodynamics in space/time split DFT [41] and cosmology [22, 42].
Moreover, it is important to note that compact spacetime is not required from the perspective of field theory. For example, in the low-energy effective action with $O\left(D,D,\mathbb{R}\right)$ duality, the spacetime coordinates are not doubled. If the background is compactified, even in field theory, the continuous $O\left(D,D,\mathbb{R}\right)$ symmetry becomes a discrete $O\left(D,D,\mathbb{Z}\right)$ symmetry. Throughout the rest of this paper, we will abbreviate $O\left(D,D,\mathbb{R}\right)$ as $O\left(D,D\right)$.
The aim of this paper is to calculate the black hole solutions of DFT. To preserve more non-trivial properties from $O\left(D,D\right)$ symmetry, we adhere to the following calculation principles: 1) Since the strong constraint erases all features of DFT, we initially calculate solutions that violate this constraint. Although these solutions currently lack clear physical interpretations, we hope future studies will solve this issue. 2) When imposing the strong constraint to obtain physical solutions, we assume the generalized metric $\mathcal{H}_{MN}$ represents the entire spacetime metric, not just the usual metric $g_{ij}$. Based on these principles, we first calculate the black hole solutions of standard DFT, which includes double time coordinates. These solutions can be derived using the following three methods. Let us recall the traditional low-energy effective action with a vanishing Kalb-Ramond field for simplicity:
$S=\int d^{4}x\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}\right].$ (1.1)
where $g_{ij}$denotes the string metric and $\phi$ is the dilaton. The black hole solutions of DFT can be constructed through the solutions of this action (1.1):
1st kind of solutions: If $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ are solutions of the effective action (1.1), the solutions of DFT that satisfy the strong constraint can be given as:
$ds^{2}=g^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{x}\right)dx^{i}dx^{j}.$ (1.2)
2nd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are self-dual solutions of the effective action (1.1), the constraint-satisfying solutions of DFT can be given as:
$ds^{2}=g^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(x\right)dx^{i}dx^{j},$ (1.3)
or
$ds^{2}=g^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(\tilde{x}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{x}\right)dx^{i}dx^{j}.$ (1.4)
3rd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are self-dual solutions of the effective action (1.1), we can construct constraint-violating solutions of DFT such that $g\left(\tilde{x},x\right)=g\left(\tilde{x}\right)\times\left(g^{-1}\right)\left(x\right)$. These solutions depend on double coordinates simultaneously and take the forms:
$ds^{2}=g^{ij}\left(x,\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x,\tilde{x}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(x,\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(x,\tilde{x}\right)dx^{i}dx^{j}.$ (1.5)
The problem with these solutions is the presence of double timelike directions $dt$ and $d\tilde{t}$, which lack explicit physical interpretations. To solve this issue, we consider the $12D-$ 1 dimensional space/time split DFT, which is independent of the dual time coordinate. To obtain non-trivial solutions, we consider non-critical strings, which introduce the cosmological constant. We then calculate the equations of motion (EOM) for the corresponding action and use the generated methods mentioned above to derive the black hole solutions. The new insights of these solutions arise from the curvature definition in doubled spacetime. In DFT, the Ricci curvature is generalized to include the doubled coordinates and dilatonic dynamics, known as the generalized Ricci scalar $\mathcal{R}$. It satisfies the EOM for the dilaton and always vanishes ($\mathcal{R}=0$). Since we consider the cosmological constant, the generalized Ricci scalar for the whole doubled spacetime is:
$\mathcal{R}=-\lambda^{2}.$ (1.6)
This implies that, regardless of the obtained solution, the entire doubled spacetime is an AdS vacuum. On the other hand, since the geometry of the subspace $x$ of the doubled spacetime can be completely determined by the low-energy effective action, we can construct the black hole solution in this subspace. In other words, considering the entire spacetime results in an AdS vacuum, whereas black holes and curvature singularities emerge in the subspace.
The remainder of this paper is outlined as follows. In Section 2, we review the dilatonic and gravitational EOM of standard DFT, classify the solutions into three kinds, and present the main examples. In Section 3, we introduce the space/time split DFT and calculate the black hole solutions of its EOM. Section 4 contains our discussion and conclusion.
2 Standard double field theory
2.1 Equations of motion for the generalized metric and dilaton
The EOM for the dilatonic and gravitational fields in DFT have been extensively reviewed in the reference [4]. Here, we briefly summarize these equations. To construct the spacetime action of DFT, we introduce the $2D\times 2D$ “generalized metric,” which incorporates the gravitational field $g_{ij}$ and the antisymmetric Kalb-Ramond field $b_{ij}$:
$\mathcal{H}_{MN}=\left(\begin{array}[]{cc}g^{ij}&-g^{ik}b_{kj}\\ b_{ik}g^{kj}&g_{ij}-b_{ik}g^{kl}b_{lj}\end{array}\right),\quad\mathcal{H}^{MN}=\left(\begin{array}[]{cc}g_{ij}-b_{ik}g^{kl}b_{lj}&b_{ik}g^{kj}\\ -g^{ik}b_{kj}&g^{ij}\end{array}\right),$ (2.7)
where $\mathcal{H}^{MP}\mathcal{H}_{PN}=\delta_{\;N}^{M}$, $\mathcal{H}^{MN}\equiv\eta^{MP}\mathcal{H}_{PQ}\eta^{QN}=\mathcal{H}_{PQ}\eta^{PM}\eta^{QN}$, and $\mathcal{H}^{MN}=\mathcal{H}^{NM}$ ($M,N,\ldots=1,2,\ldots,2D$). The indices can be raised and lowered by the $O\left(D,D\right)$-invariant metric:
$\eta_{MN}=\left(\begin{array}[]{cc}0&I\\ I&0\end{array}\right),\qquad\eta^{MN}=\eta_{MN}.$ (2.8)
The gauge-invariant DFT action can thus be constructed from $O\left(D,D\right)$ scalars:
$S$ $=$ $\int d^{D}xd^{D}\tilde{x}\mathcal{L}_{\mathrm{DFT}}\left(d,\mathcal{H}_{MN}\right)$ (2.9)
$=$ $\int d^{D}xd^{D}\tilde{x}e^{-2d}\left(\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{N}\mathcal{H}^{KL}\partial_{L}\mathcal{H}_{MK}\right.$
$\left.-2\partial_{M}d\partial_{N}\mathcal{H}^{MN}+4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d\right),$
where the derivatives are $\partial_{M}=\left(\tilde{\partial}^{i},\partial_{i}\right)$ and the $O\left(D,D\right)$-invariant dilaton $d$ is defined in terms of the physical dilaton $\phi$ as:
$e^{-2d}=\sqrt{g}e^{-2\phi}.$ (2.10)
Since each term in the action is a scalar, the $O\left(D,D\right)$ symmetry is manifest. Moreover, the action (2.9) can also be rewritten in terms of the generalized scalar curvature $\mathcal{R}\left(d,\mathcal{H}_{MN}\right)$ with some additional boundary terms:
$\mathcal{L}_{\mathrm{DFT}}\left(d,\mathcal{H}_{MN}\right)=e^{-2d}\mathcal{R}\left(d,\mathcal{H}_{MN}\right)+\partial_{M}\left(e^{-2d}\left[\partial_{N}\mathcal{H}^{MN}-4\mathcal{H}^{MN}\partial_{N}d\right]\right),$ (2.11)
where the generalized curvature scalar is defined as:
$\mathcal{R}\left(d,\mathcal{H}_{MN}\right)$ $=$ $\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{N}\mathcal{H}^{KL}\partial_{L}\mathcal{H}_{MK}+4\mathcal{H}^{MN}\partial_{M}\partial_{N}d$ (2.12)
$+4\partial_{M}\mathcal{H}^{MN}\partial_{N}d-4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d-\partial_{M}\partial_{N}\mathcal{H}^{MN}.$
Additionally, all fields must satisfy the strong constraint $\eta^{MN}\partial_{M}\partial_{N}f=0$ (or equivalently $\partial^{M}f\partial_{M}g=0$)due to the level-matching condition of closed string theory. This constraint requires all fields to live only on a subspace of the doubled coordinates. The gauge transformations of the DFT action (2.9) are given by:
$\delta_{\xi}\mathcal{H}^{MN}$ $=$ ^$\hat{\mathcal{L}}_{\xi}\mathcal{H}^{MN}\equiv\xi^{P}\partial_{P}\mathcal{H}^{MN}+\left(\partial^{M}\xi_{P}-\partial_{P}\xi^{M}\right)\mathcal{H}^{PN}+\left(\partial^{N}\xi_{P}-\partial_{P}\xi^{N}\right)\mathcal{H}^{MP},$
$\delta d$ $=$ $\xi^{M}\partial_{M}d-\frac{1}{2}\partial_{M}\xi^{M},$ (2.13)
with $\xi^{M}=\left(\hat{\xi}_{i},\xi^{i}\right)$ and “generalized Lie derivatives” ^$\hat{\mathcal{L}}_{\xi}$. The dilatonic EOM can be derived from the variation of the DFT action: (2.9):
$\delta_{d}S$ $=$ $\int dxd\tilde{x}e^{-2d}\left(-\frac{1}{4}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}+\mathcal{H}^{MN}\partial_{N}\mathcal{H}^{KL}\partial_{L}\mathcal{H}_{MK}\right.$ (2.14)
$\left.+2\partial_{M}\partial_{N}\mathcal{H}^{MN}+8\mathcal{H}^{MN}\partial_{M}d\partial_{N}d-8\partial_{M}\mathcal{H}^{MN}\partial_{N}d-8\mathcal{H}^{MN}\partial_{M}\partial_{N}d\right)\delta d.$
This gives:
$\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{K}\mathcal{H}_{NL}$
$-\partial_{M}\partial_{N}\mathcal{H}^{MN}-4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d+4\partial_{M}\mathcal{H}^{MN}\partial_{N}d+4\mathcal{H}^{MN}\partial_{M}\partial_{N}d=0.$ (2.15)
To obtain the gravitational field equation for the generalized metric $\mathcal{H}$, we consider the variation of $\mathcal{H}^{MN}$:
[…]
where
$\mathcal{K}_{MN}$$\equiv$ $\frac{1}{8}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{4}\left(\partial_{L}-2\partial_{L}d\right)\left(\mathcal{H}^{LK}\partial_{K}\mathcal{H}_{MN}\right)+2\partial_{M}\partial_{N}d$ (2.17)
$-\frac{1}{2}\partial_{\left(N\right.}\mathcal{H}^{KL}\partial_{L}\mathcal{H}_{\left.M\right)K}+\frac{1}{2}\left(\partial_{L}-2\partial_{L}d\right)\left(\mathcal{H}^{KL}\partial_{\left(N\right.}\mathcal{H}_{\left.M\right)K}+\mathcal{H}_{\quad\left(M\right.}^{K}\partial_{K}\mathcal{H}_{\quad\left.N\right)}^{L}\right).$
However, $\mathcal{K}_{MN}=0$ is not the gravitational field equation, as this equation is not an $O\left(D,D\right)$-invariant field equation. To obtain the $O\left(D,D\right)$-invariant EOM, we first utilize the notations $\mathcal{H}\equiv\mathcal{H}^{\bullet\bullet}$ and $\eta\equiv\mathcal{\eta}_{\bullet\bullet}$. Since the $O\left(D,D\right)$ symmetry requires $\mathcal{H}$ to satisfy $\mathcal{H}\eta\mathcal{H}=\eta^{-1}$, its varied field $\mathcal{H}^{\prime}=\mathcal{H}+\delta\mathcal{H}$ also needs to satisfy the symmetric condition $\mathcal{H}^{\prime}\eta\mathcal{H}^{\prime}=\eta^{-1}$. Then we have:
$\delta\mathcal{H}\eta\mathcal{H}+\mathcal{H}\eta\delta\mathcal{H}=0.$ (2.18)
To simplify our calculations, we define the following:
$S_{\quad N}^{M}\equiv\mathcal{H}_{\quad N}^{M}=\eta^{MP}\mathcal{H}_{PN}=\mathcal{H}^{MP}\eta_{PN},\quad S^{2}=1.$ (2.19)
Using the notation $S\equiv S_{\quad\bullet}^{\bullet}=\mathcal{H}\eta$, we derive the variation condition for
$\delta\mathcal{H}S^{t}+S\delta\mathcal{H}=0.$ (2.20)
Given that $S^{2}=1$, we can rewrite the variation as:
$\delta\mathcal{H}=-S\delta\mathcal{H}S^{t}.$ (2.21)
This can be further expressed as:
$\delta\mathcal{H}=\frac{1}{4}\left(1+S\right)\mathcal{M}\left(1-S^{t}\right)+\frac{1}{4}\left(1-S\right)\mathcal{M}\left(1+S^{t}\right),$ (2.22)
where $\mathcal{M}$ is an arbitrary symmetric matrix ensuring the symmetry of $\delta\mathcal{H}$. Substituting this back into the variation of the action (2.16), we get:
$\frac{1}{4}\left(1-S^{t}\right)\mathcal{K}\left(1+S\right)+\frac{1}{4}\left(1+S^{t}\right)\mathcal{K}\left(1-S\right)=0.$ (2.23)
Thus, the $O\left(D,D\right)$-invariant gravitational field equation is obtained:
$\mathcal{R}_{MN}$ $=$ $\frac{1}{4}\left(\delta_{M}^{\quad P}-S_{\quad M}^{P}\right)\mathcal{K}_{PQ}\left(\delta_{\quad N}^{Q}+S_{\quad N}^{Q}\right)+\frac{1}{4}\left(\delta_{M}^{\quad P}+S_{\quad M}^{P}\right)\mathcal{K}_{PQ}\left(\delta_{\quad N}^{Q}-S_{\quad N}^{Q}\right)$ (2.24)
$=$ $\frac{1}{2}\mathcal{K}_{MN}-\frac{1}{2}S_{\quad M}^{P}\mathcal{K}_{PQ}S_{\quad N}^{Q}.$
In summary, the dilatonic and gravitational EOM of DFT are presented as follows:
$\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{K}\mathcal{H}_{NL}-\partial_{M}\partial_{N}\mathcal{H}^{MN}$
$-4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d+4\partial_{M}\mathcal{H}^{MN}\partial_{N}d+4\mathcal{H}^{MN}\partial_{M}\partial_{N}d$ $=$ $0,$
$\mathcal{K}_{MN}-S_{\quad M}^{P}\mathcal{K}_{PQ}S_{\quad N}^{Q}$ $=$ $0$ (2.25)
2.2 Some notations and definitions
To calculate the EOM (2.25) in DFT, we need to define several notations and calculation rules. First, we clarify the notations for the generalized metric. The spatial vector in DFT is defined as:
$\partial_{M}\equiv\left(\begin{array}[]{c}\tilde{\partial}^{i}\\ \partial_{i}\end{array}\right),$ (2.26)
where $i$ denotes the $D-1$ dimensional spatial coordinates. The corresponding dual vector is given by:
$dX^{M}\equiv\left(\begin{array}[]{c}d\tilde{x}_{i}\\ dx^{i}\end{array}\right).$ (2.27)
Using the dual vector, we can define the generalized line element in terms of the generalized metric:
$ds^{2}=\mathcal{H}_{MN}dX^{M}dX^{N}.$ (2.28)
For simplicity, we use block matrix notation to rewrite the vector, the dual vector, and the generalized metric:
$\partial_{M}=\left(\begin{array}[]{c}\partial_{\mathbf{1}}\\ \partial_{\mathbf{2}}\end{array}\right),\quad dX^{M}=\left(\begin{array}[]{c}dX^{\mathbf{\mathbf{1}}}\\ dX^{\mathbf{2}}\end{array}\right),\quad\mathcal{H}_{MN}=\left(\begin{array}[]{cc}\mathcal{H}_{\mathbf{1}\mathbf{1}}&\mathcal{H}_{\mathbf{1}\mathbf{2}}\\ \mathcal{H}_{\mathbf{2}\mathbf{1}}&\mathcal{H}_{\mathbf{2}\mathbf{2}}\end{array}\right),\quad\mathcal{H}^{MN}=\left(\begin{array}[]{cc}\mathcal{H}^{\mathbf{1}\mathbf{1}}&\mathcal{H}^{\mathbf{1}\mathbf{2}}\\ \mathcal{H}^{\mathbf{2}\mathbf{1}}&\mathcal{H}^{\mathbf{2}\mathbf{2}}\end{array}\right),$ (2.29)
where $\mathbf{1}$ represents the set of coordinates $\tilde{x}_{i}$ and $\mathbf{2}$ denotes the coordinates $x^{i}$.Using these notations, the components of the generalized metric can be divided into four blocks:
$\mathcal{H}_{\mathbf{1}\mathbf{1}}\left(\frac{\partial}{\partial\tilde{x}_{i}},\frac{\partial}{\partial\tilde{x}_{j}}\right)=g^{ij},$ $\mathcal{H}_{\mathbf{1}\mathbf{2}}\left(\frac{\partial}{\partial\tilde{x}_{i}},\frac{\partial}{\partial x^{j}}\right)=-g^{ik}b_{kj},$
$\mathcal{H}_{\mathbf{2}\mathbf{1}}\left(\frac{\partial}{\partial x^{i}},\frac{\partial}{\partial\tilde{x}_{j}}\right)=b_{ik}g^{kj},$ $\mathcal{H}_{\mathbf{2}\mathbf{2}}\left(\frac{\partial}{\partial x^{i}},\frac{\partial}{\partial x^{j}}\right)=g_{ij}-b_{ik}g^{kl}b_{lj}.$ (2.30)
It is worth noting that $g\left(x,\tilde{x}\right)$, $b\left(x,\tilde{x}\right)$, and the dilaton $d\left(x,\tilde{x}\right)$ depend on two sets of coordinates. The generalized line element becomes:
$dS^{2}$ $=$ $\mathcal{H}_{\mathbf{1}\mathbf{1}}dX^{\mathbf{1}}dX^{\mathbf{1}}+\mathcal{H}_{\mathbf{1}\mathbf{2}}dX^{\mathbf{1}}dX^{\mathbf{2}}+\mathcal{H}_{\mathbf{2}\mathbf{1}}dX^{\mathbf{2}}dX^{\mathbf{1}}+\mathcal{H}_{\mathbf{2}\mathbf{2}}dX^{\mathbf{2}}dX^{\mathbf{2}}$ (2.31)
$=$ $g^{ij}d\tilde{x}_{i}d\tilde{x}_{j}-g^{ik}b_{kj}d\tilde{x}_{i}dx^{j}+b_{ik}g^{kj}dx^{i}d\tilde{x}_{j}+\left(g_{ij}-b_{ik}g^{kl}b_{lj}\right)dx^{i}dx^{j}.$
For simplicity in the rest of the paper, we will set the Kalb-Ramond field to be zero, so the line element becomes:
$dS^{2}=\mathcal{H}_{\mathbf{1}\mathbf{1}}dX^{\mathbf{1}}dX^{\mathbf{1}}+\mathcal{H}_{\mathbf{2}\mathbf{2}}dX^{\mathbf{2}}dX^{\mathbf{2}}=g^{ij}d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}dx^{i}dx^{j}.$ (2.32)
Then, the blocks of the generalized metric are:
$\mathcal{H}_{\mathbf{1}\mathbf{1}}=g^{ij},\qquad\mathcal{H}^{\mathbf{1}\mathbf{1}}=g_{ij},\qquad\mathcal{H}^{\mathbf{2}\mathbf{2}}=g^{ij},\qquad\mathcal{H}_{\mathbf{\mathbf{2}}\mathbf{2}}=g_{ij}.$ (2.33)
To calculate the contraction of $O\left(D,D\right)$ indices, we introduce extra indices $\mathcal{H}^{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}$ to represent the contractions of the elements of the block matrix. Therefore, the generalized metric can be rewritten by adding the extra indices:
$\mathcal{H}_{M\left(i\right)N\left(j\right)}=\left(\begin{array}[]{cc}\mathcal{H}_{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}&0\\ 0&\mathcal{H}_{\mathbf{2}\left(i\right)\mathbf{2}\left(j\right)}\end{array}\right)=\left(\begin{array}[]{cc}g^{ij}&0\\ 0&g_{ij}\end{array}\right),$ (2.34)
where $b_{ij}=0$. Now, there exist two sets of indices: the block matrix notations $M$, $N$ and the indices of the components of the block matrix $i$, $j$. Based on these new notations, we give the following examples that we will use in the calculations:
$\mathcal{H}^{M\left(i\right)N\left(j\right)}\partial_{M\left(i\right)}d\partial_{N\left(j\right)}d$ $=$ $\mathcal{H}^{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}\partial_{\mathbf{1}\left(i\right)}d\partial_{\mathbf{1}\left(j\right)}d+\mathcal{H}^{\mathbf{2}\left(i\right)\mathbf{2}\left(j\right)}\partial_{\mathbf{2}\left(i\right)}d\partial_{\mathbf{2}\left(j\right)}d$ (2.35)
$=$ $g_{ij}\tilde{\partial}^{i}d\tilde{\partial}^{j}d+g^{ij}\partial_{i}d\partial_{j}d.$
Moreover, for the derivatives, we have:
$\partial_{\mathbf{1}\left(k\right)}\mathcal{H}^{\mathbf{2}\left(i\right)\mathbf{2}\left(j\right)}$ $=$ ~$\tilde{\partial}^{k}g^{ij}$
$\partial_{\mathbf{2}\left(k\right)}\mathcal{H}^{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}$ $=$ $\partial_{k}g_{ij}$
$\partial_{\mathbf{1}\left(p\right)}\mathcal{H}^{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}\partial_{\mathbf{2}\left(q\right)}\mathcal{H}_{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}$ $=$ ~$\tilde{\partial}^{p}g_{ij}\partial_{q}g^{ij}$ (2.36)
On the other hand, the rules for $S_{\quad N}^{M}$ are
$\mathcal{H}_{\quad\mathbf{1}\left(j\right)}^{\mathbf{2}\left(i\right)}=g^{ij},\quad\mathcal{H}_{\quad\mathbf{2}\left(j\right)}^{\mathbf{1}\left(i\right)}=g_{ij}.$ (2.37)
and
$\partial_{\mathbf{2}\left(k\right)}\mathcal{H}_{\quad\mathbf{2}\left(j\right)}^{\mathbf{1}\left(i\right)}=\partial_{k}g_{ij},\qquad\partial_{\mathbf{1}\left(k\right)}\mathcal{H}_{\quad\mathbf{1}\left(j\right)}^{\mathbf{2}\left(i\right)}=\tilde{\partial}^{k}g^{ij}.$ (2.38)
Note that the raising and lowering of indices are given by:
$g_{ij}\tilde{\partial}^{i}=\tilde{\partial}_{j},\quad g^{ij}\partial_{i}=\partial^{j}.$ (2.39)
2.3 Classification of black hole solutions
In this subsection, our goal is to outline the methodology for generating typical DFT black hole solutions. These solutions are categorized into three types, each accompanied by specific examples. However, due to the inclusion of dual time coordinates, the physical interpretation of these solutions is not explicitly defined.
Firstly, let us revisit the EOM for the graviton and dilaton fields:
$\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{K}\mathcal{H}_{NL}-\partial_{M}\partial_{N}\mathcal{H}^{MN}$
$-4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d+4\partial_{M}\mathcal{H}^{MN}\partial_{N}d+4\mathcal{H}^{MN}\partial_{M}\partial_{N}d$ $=$ $0,$
$2\mathcal{R}_{MN}=\mathcal{K}_{MN}-S_{\quad M}^{P}\mathcal{K}_{PQ}S_{\quad N}^{Q}$ $=$ $0$ (2.40)
Assuming the Kalb-Ramond field is zero and substituting the generalized metric formulation:
$\mathcal{H}_{M\left(i\right)N\left(j\right)}=\left(\begin{array}[]{cc}\mathcal{H}_{\mathbf{1}\left(i\right)\mathbf{1}\left(j\right)}&0\\ 0&\mathcal{H}_{\mathbf{2}\left(i\right)\mathbf{2}\left(j\right)}\end{array}\right)=\left(\begin{array}[]{cc}g^{ij}&0\\ 0&g_{ij}\end{array}\right),$ (2.41)
the EOM can be expanded using the notation and rules established in the previous section. Specifically for the graviton field, we obtain:
$\mathcal{R}_{\mathbf{1}\left(p\right)\mathbf{1}\left(q\right)}$ $=$ $\frac{1}{8}\tilde{\partial}^{p}g_{ij}\tilde{\partial}^{q}g^{ij}+\frac{1}{8}\tilde{\partial}^{p}g^{ij}\tilde{\partial}^{q}g_{ij}-\frac{1}{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}g^{pq}\right)-\frac{1}{4}\partial_{i}\left(g^{ij}\partial_{j}g^{pq}\right)$
$-\frac{1}{4}\tilde{\partial}^{q}g_{ij}\tilde{\partial}^{j}g^{pi}-\frac{1}{4}\tilde{\partial}^{p}g_{ij}\tilde{\partial}^{j}g^{qi}+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\tilde{\partial}^{q}g^{pj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\tilde{\partial}^{p}g^{qj}\right)$
$+\frac{1}{4}\partial_{i}\left(g^{jp}\partial_{j}g^{iq}\right)+\frac{1}{4}\partial_{i}\left(g^{jq}\partial_{j}g^{ip}\right)-\frac{1}{8}g^{ip}\partial_{i}g_{kl}\partial_{j}g^{kl}g^{jq}-\frac{1}{8}g^{ip}\partial_{i}g^{kl}\partial_{j}g_{kl}g^{jq}$
$+\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{lk}\tilde{\partial}^{k}g_{ij}\right)g^{jq}+\frac{1}{4}g^{ip}\partial_{l}\left(g^{lk}\partial_{k}g_{ij}\right)g^{jq}+\frac{1}{4}g^{ip}\partial_{j}g^{kl}\partial_{l}g_{ik}g^{jq}+\frac{1}{4}g^{ip}\partial_{i}g^{kl}\partial_{l}g_{jk}g^{jq}$
$-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\partial_{j}g_{ik}\right)g^{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\partial_{i}g_{jk}\right)g^{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{ki}\tilde{\partial}^{k}g_{lj}\right)g^{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{kj}\tilde{\partial}^{k}g_{li}\right)g^{jq}$
$+\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}g^{pq}\right)+\frac{1}{2}\partial_{i}d\left(g^{ij}\partial_{j}g^{pq}\right)+2\tilde{\partial}^{p}\tilde{\partial}^{q}d$
$-\frac{1}{2}\partial_{i}d\left(g^{jp}\partial_{j}g^{iq}\right)-\frac{1}{2}\partial_{i}d\left(g^{jq}\partial_{j}g^{ip}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ji}\tilde{\partial}^{q}g^{pj}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ji}\tilde{\partial}^{p}g^{qj}\right)$
$-\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{lk}\tilde{\partial}^{k}g_{ij}\right)g^{jq}-\frac{1}{2}g^{ip}\partial_{l}d\left(g^{lk}\partial_{k}g_{ij}\right)g^{jq}-2g^{ip}\partial_{i}\partial_{j}dg^{jq}$
$+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\partial_{j}g_{ik}\right)g^{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\partial_{i}g_{jk}\right)g^{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{ki}\tilde{\partial}^{k}g_{lj}\right)g^{jq}+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{kj}\tilde{\partial}^{k}g_{li}\right)g^{jq}.$
$\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{2}\left(q\right)}$ $=$ $\frac{1}{8}\partial_{p}g_{ij}\partial_{q}g^{ij}+\frac{1}{8}\partial_{p}g^{ij}\partial_{q}g_{ij}-\frac{1}{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}g_{pq}\right)-\frac{1}{4}\partial_{i}\left(g^{ij}\partial_{j}g_{pq}\right)$
$-\frac{1}{4}\partial_{q}g^{ij}\partial_{j}g_{pi}-\frac{1}{4}\partial_{p}g^{ij}\partial_{j}g_{qi}+\frac{1}{4}\partial_{i}\left(g^{ji}\partial_{q}g_{pj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\partial_{p}g_{qj}\right)$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{jp}\tilde{\partial}^{j}g_{iq}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g_{jq}\tilde{\partial}^{j}g_{ip}\right)-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{j}g^{kl}g_{jq}-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g^{kl}\tilde{\partial}^{j}g_{kl}g_{jq}$
$+\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{lk}\tilde{\partial}^{k}g^{ij}\right)g_{jq}+\frac{1}{4}g_{ip}\partial_{l}\left(g^{lk}\partial_{k}g^{ij}\right)g_{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{j}g_{kl}\tilde{\partial}^{l}g^{ik}g_{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{l}g^{jk}g_{jq}$
$-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\tilde{\partial}^{j}g^{ik}\right)g_{jq}-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\tilde{\partial}^{i}g\right)g_{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{ki}\partial_{k}g^{lj}\right)g_{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{kj}\partial_{k}g^{li}\right)g_{jq}$
$+\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}g_{pq}\right)+\frac{1}{2}\partial_{i}d\left(g^{ij}\partial_{j}g_{pq}\right)+2\partial_{p}\partial_{q}d$
$-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{jp}\tilde{\partial}^{j}g_{iq}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{jq}\tilde{\partial}^{j}g_{ip}\right)-\frac{1}{2}\partial_{i}d\left(g^{ji}\partial_{q}g_{pj}\right)-\frac{1}{2}\partial_{i}d\left(g^{ji}\partial_{p}g_{qj}\right)$
$-\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{lk}\tilde{\partial}^{k}\tilde{g}^{ij}\right)g_{jq}-\frac{1}{2}g_{ip}\partial_{l}d\left(g^{lk}\partial_{k}g^{ij}\right)g_{jq}-2g_{ip}\tilde{\partial}^{i}\tilde{\partial}^{j}dg_{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\tilde{\partial}^{j}g^{ik}\right)g_{jq}+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\tilde{\partial}^{i}g^{jk}\right)g_{jq}$
$+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{ki}\partial_{k}g^{lj}\right)g_{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{kj}\partial_{k}g^{li}\right)g_{jq},$
$\mathcal{R}_{\mathbf{1}\left(p\right)\mathbf{2}\left(q\right)}$ $=$ $\frac{1}{8}\tilde{\partial}^{p}g_{ij}\partial_{q}g^{ij}+\frac{1}{8}\tilde{\partial}^{p}g^{ij}\partial_{q}g_{ij}-\frac{1}{4}\partial_{q}g_{ij}\tilde{\partial}^{j}g^{pi}-\frac{1}{4}\tilde{\partial}^{p}g^{ij}\partial_{j}g_{qi}$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\partial_{q}g^{pj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\tilde{\partial}^{p}g_{qj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g^{jp}\partial_{j}g_{iq}\right)+\frac{1}{4}\partial_{i}\left(g_{jq}\tilde{\partial}^{j}g^{ip}\right)$
$-\frac{1}{8}g^{ip}\partial_{i}g_{kl}\tilde{\partial}^{j}g^{kl}g_{jq}-\frac{1}{8}g^{ip}\partial_{i}g^{kl}\tilde{\partial}^{j}g_{kl}g_{jq}+\frac{1}{4}g^{ip}\partial_{i}g_{kl}\tilde{\partial}^{l}g^{jk}g_{jq}+\frac{1}{4}g^{ip}\tilde{\partial}^{j}g^{kl}\partial_{l}g_{ik}g_{jq}$
$-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{kl}\partial_{i}g^{jk}\right)g_{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\tilde{\partial}^{j}g_{ik}\right)g_{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g_{ki}\tilde{\partial}^{k}g^{lj}\right)g_{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g^{kj}\partial_{k}g_{li}\right)g_{jq}$
$+2\tilde{\partial}^{p}\partial_{q}d-\tilde{\partial}^{i}d\left(g^{jp}\partial_{j}g_{iq}\right)-\partial_{i}d\left(g_{jq}\tilde{\partial}^{j}g^{ip}\right)$
~$-\tilde{\partial}^{i}d\left(g_{ji}\partial_{q}g^{pj}\right)-\partial_{i}d\left(g^{ji}\tilde{\partial}^{p}g_{qj}\right)-2g^{ip}\partial_{i}\tilde{\partial}^{j}dg_{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{kl}\partial_{i}g^{jk}\right)g_{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\tilde{\partial}^{j}g_{ik}\right)g_{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g^{kj}\partial_{k}g_{li}\right)g_{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g_{ki}\tilde{\partial}^{k}g^{lj}\right)g_{jq},$
$\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{1}\left(q\right)}$ $=$ $\frac{1}{8}\partial_{p}g_{ij}\tilde{\partial}^{q}g^{ij}+\frac{1}{8}\partial_{p}g^{ij}\tilde{\partial}^{q}g_{ij}-\frac{1}{4}\partial_{p}g_{ij}\tilde{\partial}^{j}g^{qi}-\frac{1}{4}\tilde{\partial}^{q}g^{ij}\partial_{j}g_{pi}$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\partial_{p}g^{qj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\tilde{\partial}^{q}g_{pj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g^{jq}\partial_{j}g_{ip}\right)+\frac{1}{4}\partial_{i}\left(g_{jp}\tilde{\partial}^{j}g^{iq}\right)$
$-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g_{kl}\partial_{j}g^{kl}g^{jq}-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g^{kl}\partial_{j}g_{kl}g^{jq}+\frac{1}{4}g_{ip}\partial_{j}g_{kl}\tilde{\partial}^{l}g^{ik}g^{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{i}g^{kl}\partial_{l}g_{jk}g^{jq}$
$-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\partial_{j}g^{ik}\right)g^{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{kl}\tilde{\partial}^{i}g_{jk}\right)g^{jq}-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g^{ki}\partial_{k}g_{lj}\right)g^{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g_{kj}\tilde{\partial}^{k}g^{li}\right)g^{jq}$
$2\partial_{p}\tilde{\partial}^{q}d-\tilde{\partial}^{i}d\left(g^{jq}\partial_{j}g_{ip}\right)-\partial_{i}d\left(g_{jp}\tilde{\partial}^{j}g^{iq}\right)$
~$-\tilde{\partial}^{i}d\left(g_{ji}\partial_{p}g^{qj}\right)-\partial_{i}d\left(g^{ji}\tilde{\partial}^{q}g_{pj}\right)-2g_{ip}\tilde{\partial}^{i}\partial_{j}dg^{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\partial_{j}g^{ik}\right)g^{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{kl}\tilde{\partial}^{i}g_{jk}\right)g^{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g^{ki}\partial_{k}g_{lj}\right)g^{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g_{kj}\tilde{\partial}^{k}g^{li}\right)g^{jq}.$
On the other hand, the EOM for the dilaton is given by:
$\frac{1}{8}g_{ij}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{j}g^{kl}+\frac{1}{8}g_{ij}\tilde{\partial}^{i}g^{kl}\tilde{\partial}^{j}g_{kl}+\frac{1}{8}g^{ij}\partial_{i}g_{kl}\partial_{j}g^{kl}+\frac{1}{8}g^{ij}\partial_{i}g^{kl}\partial_{j}g_{kl}$
$-\frac{1}{2}g_{ij}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{k}g^{jl}-\frac{1}{2}g^{ij}\partial_{i}g^{kl}\partial_{k}g_{jl}-\tilde{\partial}^{i}\tilde{\partial}^{j}g_{ij}-\partial_{i}\partial_{j}g^{ij}-4g_{ij}\tilde{\partial}^{i}d\tilde{\partial}^{j}d-4g^{ij}\partial_{i}d\partial_{j}d$
$+4\tilde{\partial}^{i}g_{ij}\tilde{\partial}^{j}d+4\partial_{i}g^{ij}\partial_{j}d+4g_{ij}\tilde{\partial}^{i}\tilde{\partial}^{j}d+4g^{ij}\partial_{i}\partial_{j}d=0.$ (2.46)
Given the symmetries between the generalized Ricci scalars $\mathcal{R}_{M\left(p\right)N\left(q\right)}$:
$\mathcal{R}_{\mathbf{1}\left(p\right)\mathbf{1}\left(q\right)}$ $\underleftrightarrow{g^{\bullet\bullet}\leftrightarrow g_{\bullet\bullet},\quad\tilde{\partial}^{\bullet}\leftrightarrow\partial_{\bullet}}$ $\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{2}\left(q\right)},$
$\mathcal{R}_{\mathbf{1}\left(p\right)\mathbf{2}\left(q\right)}$ $\underleftrightarrow{g^{\bullet\bullet}\leftrightarrow g_{\bullet\bullet},\quad\tilde{\partial}^{\bullet}\leftrightarrow\partial_{\bullet}}$ $\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{1}\left(q\right)},$ (2.47)
we need only compute one set. In this paper, we focus on calculating $\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{2}\left(q\right)}$ and $\mathcal{R}_{\mathbf{2}\left(p\right)\mathbf{1}\left(q\right)}$.
Next, we present three types of black hole solutions for the EOM (LABEL:eq:_GRA_EOM) and (2.46). To begin, we revisit the low-energy effective action of closed string theory:
$S=\int d^{4}x\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}-\frac{1}{12}H_{\mu\nu\rho}H^{\mu\nu\rho}\right].$ (2.48)
The solutions of standard DFT black holes are categorized by the solutions $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ from the action (2.48), as follows:
1st kind of black hole solutions
If $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ satisfy the action (2.48), they automatically become solutions of DFT that satisfy the strong constraint. The corresponding line elements can take two forms:
$ds^{2}=g^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{x}\right)dx^{i}dx^{j}.$ (2.49)
Moreover, when $\phi\left(x\right)=0$, the low-energy effective action (2.48) reduces to the Einstein-Hilbert action, implying that DFT also includes the black hole solutions of Einstein’s gravity. There is a notable exception: the BTZ black hole. The BTZ black hole cannot be directly obtained by setting $\phi\left(x\right)=0$; it also requires the inclusion of the Kalb-Ramond field ($H_{ijk}\neq 0$) to precisely reduce to Einstein’s equations with a negative cosmological constant [43]. Therefore, it is necessary to consider a non-vanishing $b_{ij}$ in this case.
Example:
To illustrate, consider the following ansatz:
$dS^{2}$ $=$ $-A\left(\tilde{r},r\right)^{-1}d\tilde{t}^{2}+B\left(\tilde{r},r\right)^{-1}d\tilde{r}^{2}+C\left(\tilde{r},r\right)^{-1}\left(d\tilde{\theta}^{2}+D\left(\tilde{\theta},\theta\right)^{-1}d\tilde{\phi}^{2}\right)$ (2.50)
$-A\left(\tilde{r},r\right)dt^{2}+B\left(\tilde{r},r\right)dr^{2}+C\left(\tilde{r},r\right)\left(d\theta^{2}+D\left(\tilde{\theta},\theta\right)d\phi^{2}\right).$
Substituting this ansatz into the EOM (LABEL:eq:_GRA_EOM) and (2.46), we obtain the EOM for the graviton:
$B\left[\left(\frac{\tilde{\partial}^{r}B}{B}-\frac{\tilde{\partial}^{r}A}{A}\right)\frac{\tilde{\partial}^{r}A}{A}+\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}A}{A}-2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}A}{A}\right]-\frac{1}{B}\left[\left(\frac{\partial_{r}B}{B}+\frac{\partial_{r}A}{A}\right)\frac{\partial_{r}A}{A}-\frac{\partial_{r}\partial_{r}A}{A}+2\partial_{r}d\frac{\partial_{r}A}{A}\right]$ $=$ $0,$
$B\left[\frac{1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}-\frac{1}{2}\left(\frac{\tilde{\partial}^{r}B}{B}\right)^{2}+\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}B}{B}-2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}B}{B}-4\tilde{\partial}^{r}\tilde{\partial}^{r}d\right]$
$-\frac{1}{B}\left[\frac{1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}+\frac{3}{2}\left(\frac{\partial_{r}B}{B}\right)^{2}+\left(\frac{\partial_{r}C}{C}\right)^{2}-\frac{\partial_{r}\partial_{r}B}{B}+2\partial_{r}d\frac{\partial_{r}B}{B}-4\partial_{r}\partial_{r}d\right]$ $=$ $0,$
$\frac{1}{2}B\left[-\frac{\tilde{\partial}^{r}B}{B}\frac{\tilde{\partial}^{r}C}{C}-\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}C}{C}+\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}C}{C}\right]-C\left[2\tilde{\partial}^{\theta}\tilde{\partial}^{\theta}d-\frac{1}{4}\left(\frac{\tilde{\partial}^{\theta}D}{D}\right)^{2}\right]$
$-\frac{1}{2}\frac{1}{B}\left[-\frac{\partial_{r}B}{B}\frac{\partial_{r}C}{C}+\frac{\partial_{r}\partial_{r}C}{C}-\left(\frac{\partial_{r}C}{C}\right)^{2}-2\partial_{r}d\frac{\partial_{r}C}{C}\right]-\frac{1}{C}\left[-2\partial_{\theta}\partial_{\theta}d+\frac{1}{4}\left(\frac{\partial_{\theta}D}{D}\right)^{2}\right]$ $=$ $0,$
$\frac{1}{2}B\left[-\frac{\tilde{\partial}^{r}B}{B}\frac{\tilde{\partial}^{r}C}{C}-\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}C}{C}+\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}C}{C}\right]-C\left[\frac{1}{2}\frac{\tilde{\partial}^{\theta}\tilde{\partial}^{\theta}D}{D}-\frac{1}{2}\left(\frac{\tilde{\partial}^{\theta}D}{D}\right)^{2}-\tilde{\partial}^{\theta}d\frac{\tilde{\partial}^{\theta}D}{D}\right]$
$-\frac{1}{2}\frac{1}{B}\left[-\frac{\partial_{r}B}{B}\frac{\partial_{r}C}{C}+\frac{\partial_{r}\partial_{r}C}{C}-\left(\frac{\partial_{r}C}{C}\right)^{2}-2\partial_{r}d\frac{\partial_{r}C}{C}\right]-\frac{1}{C}\left[\frac{1}{2}\frac{\partial_{\theta}\partial_{\theta}D}{D}-\frac{1}{2}\left(\frac{\partial_{\theta}D}{D}\right)^{2}-\partial_{\theta}d\frac{\partial_{\theta}D}{D}\right]$ $=$ $0,$
and the EOM for the dilaton:
$\frac{1}{2}B\left(-\frac{1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}+\frac{1}{2}\left(\frac{\tilde{\partial}^{r}B}{B}\right)^{2}-\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}-2\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}B}{B}+8\frac{\tilde{\partial}^{r}B}{B}\tilde{\partial}^{r}d+8\tilde{\partial}^{r}\tilde{\partial}^{r}d-8\tilde{\partial}^{r}d\tilde{\partial}^{r}d\right)$
$+C\left[4\tilde{\partial}^{\theta}\tilde{\partial}^{\theta}d-4\tilde{\partial}^{\theta}d\tilde{\partial}^{\theta}d-\frac{1}{4}\left(\frac{\tilde{\partial}^{\theta}D}{D}\right)^{2}\right]$
$+\frac{1}{2}\frac{1}{B}\left(-\frac{1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}-\frac{7}{2}\left(\frac{\partial_{r}B}{B}\right)^{2}-\left(\frac{\partial_{r}C}{C}\right)^{2}+2\frac{\partial_{r}\partial_{r}B}{B}-8\frac{\partial_{r}B}{B}\partial_{r}d+8\partial_{r}\partial_{r}d-8\partial_{r}d\partial_{r}d\right)$
$+\frac{1}{C}\left[4\partial_{\theta}\partial_{\theta}d-4\partial_{\theta}d\partial_{\theta}d-\frac{1}{4}\left(\frac{\partial_{\theta}D}{D}\right)^{2}\right]=0.$ (2.52)
Keep in mind that ~$\tilde{\partial}^{r}\equiv\frac{\partial}{\partial\tilde{r}}$ and ~$\tilde{\partial}^{\theta}\equiv\frac{\partial}{\partial\tilde{\theta}}$. After imposing the strong constraint, the solutions are organized as follows. The first solution is:
$dS^{2}$ $=$ $-A\left(r\right)^{-1}d\tilde{t}^{2}+B\left(r\right)^{-1}d\tilde{r}^{2}+C\left(r\right)^{-1}\left(d\tilde{\theta}^{2}+D\left(\theta\right)^{-1}d\tilde{\phi}^{2}\right)$ (2.53)
$-A\left(r\right)dt^{2}+B\left(r\right)dr^{2}+C\left(r\right)\left(d\theta^{2}+D\left(\theta\right)d\phi^{2}\right),$
where
$A\left(r\right)$ $=$ $\left(1-\frac{2\eta}{r}\right)^{\frac{m+\sigma}{\eta}},$
$B\left(r\right)$ $=$ $\left(1-\frac{2\eta}{r}\right)^{\frac{\sigma-m}{\eta}},$
$C\left(r\right)$ $=$ $\left(1-\frac{2\eta}{r}\right)^{1+\frac{\sigma-m}{\eta}}r^{2},$
$D\left(\theta\right)$ $=$ $\sin^{2}\theta,$
$d\left(r,\theta\right)$ $=$ $\frac{\sigma}{2\eta}\ln\left(1-\frac{2\eta}{r}\right)-\frac{1}{4}\log\left[r^{4}\left(1-\frac{2\eta}{r}\right)^{2+3\frac{\sigma-m}{\eta}+\frac{m+\sigma}{\eta}}\sin^{2}\theta\right].$ (2.54)
Its dual solution can be obtained similarly using the following ansatz:
$dS^{2}$ $=$ $-A\left(\tilde{r}\right)d\tilde{t}^{2}+B\left(\tilde{r}\right)d\tilde{r}^{2}+C\left(\tilde{r}\right)\left(d\tilde{\theta}^{2}+D\left(\tilde{\theta}\right)d\tilde{\phi}^{2}\right)$ (2.55)
$-A\left(\tilde{r}\right)^{-1}dt^{2}+B\left(\tilde{r}\right)^{-1}dr^{2}+C\left(\tilde{r}\right)^{-1}\left(d\theta^{2}+D\left(\tilde{\theta}\right)^{-1}d\phi^{2}\right),$
and its solution is
$A\left(\tilde{r}\right)$ $=$ $\left(1-\frac{2\eta}{\tilde{r}}\right)^{\frac{m+\sigma}{\eta}},$
$B\left(\tilde{r}\right)$ $=$ $\left(1-\frac{2\eta}{\tilde{r}}\right)^{\frac{\sigma-m}{\eta}},$
$C\left(\tilde{r}\right)$ $=$ $\left(1-\frac{2\eta}{\tilde{r}}\right)^{1+\frac{\sigma-m}{\eta}}\tilde{r}^{2},$
$D\left(\tilde{\theta}\right)$ $=$ $\sin^{2}\tilde{\theta},$
$d\left(\tilde{r},\tilde{\theta}\right)$ $=$ $\frac{\sigma}{2\eta}\ln\left(1-\frac{2\eta}{\tilde{r}}\right)-\frac{1}{4}\log\left[\tilde{r}^{4}\left(1-\frac{2\eta}{\tilde{r}}\right)^{2+3\frac{\sigma-m}{\eta}+\frac{m+\sigma}{\eta}}\sin^{2}\tilde{\theta}\right],$ (2.56)
where $m$ is the mass of the black hole, $\sigma$ is the scalar charge, and $\eta^{2}=m^{2}+\sigma^{2}$. When $\sigma=0$, these reduce to doubly Schwarzschild black holes. A similar solution in the low-energy effective action was found in [44].
2nd kind of black hole solutions
If $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ are solutions of the action (2.48), and $\left(g^{-1}\right)_{ij}\left(x\right)$ is also a solution for the same action, these configurations are termed self-dual solutions. They satisfy the strong constraint and encompass four distinct types, expressed as:
$ds^{2}=g^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(x\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(x\right)dx^{i}dx^{j}{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25},}$ (2.57)
or
$ds^{2}=g^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(\tilde{x}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{x}\right)dx^{i}dx^{j}.$ (2.58)
Example:
To illustrate this self-dual property, consider the following ansatz:
$dS^{2}$ $=$ $-A\left(\tilde{r},r\right)^{-1}d\tilde{t}^{2}+d\tilde{r}^{2}+A\left(\tilde{r},r\right)^{-1}\left(d\tilde{x}_{i}d\tilde{x}^{i}\right)$ (2.59)
$-A\left(\tilde{r},r\right)dt^{2}+dr^{2}+A\left(\tilde{r},r\right)\left(dx_{i}dx^{i}\right).$
The EOM (LABEL:eq:_GRA_EOM) and (2.46) for the graviton are then:
[(~$\left[-\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}+\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}A}{A}-2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}A}{A}\right]-\left[\left(\frac{\partial_{r}A}{A}\right)^{2}-\frac{\partial_{r}\partial_{r}A}{A}+2\partial_{r}d\frac{\partial_{r}A}{A}\right]$ $=$ $0,$
$\left[\frac{D-1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}-4\tilde{\partial}^{r}\tilde{\partial}^{r}d\right]-\left[\frac{D-1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}-4\partial_{r}\partial_{r}d\right]$ $=$ $0,$ (2.60)
while the EOM for the dilaton is:
$\left(-\frac{D-1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}+8\tilde{\partial}^{r}\tilde{\partial}^{r}d-8\tilde{\partial}^{r}d\tilde{\partial}^{r}d\right)+\left(-\frac{D-1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}+8\partial_{r}\partial_{r}d-8\partial_{r}d\partial_{r}d\right)=0.$ (2.61)
This formulation yields two types of solutions, given by:
$dS^{2}=-\left(\frac{r}{r_{0}}\right)^{\frac{\pm 2}{\sqrt{D-1}}}d\tilde{t}^{2}+d\tilde{r}^{2}+\left(\frac{r}{r_{0}}\right)^{\frac{\pm 2}{\sqrt{D-1}}}\left(d\tilde{x}_{i}d\tilde{x}^{i}\right)-\left(\frac{r}{r_{0}}\right)^{\frac{\mp 2}{\sqrt{D-1}}}dt^{2}+dr^{2}+\left(\frac{r}{r_{0}}\right)^{\frac{\mp 2}{\sqrt{D-1}}}\left(dx_{i}dx^{i}\right),$ (2.62)
which depend on the usual coordinate $r$, and two additional types of solutions that depend on the dual coordinate $\tilde{r}$:
$dS^{2}=-\left(\frac{\tilde{r}}{\tilde{r_{0}}}\right)^{\frac{\pm 2}{\sqrt{D-1}}}d\tilde{t}^{2}+d\tilde{r}^{2}+\left(\frac{\tilde{r}}{\tilde{r_{0}}}\right)^{\frac{\pm 2}{\sqrt{D-1}}}\left(d\tilde{x}_{i}d\tilde{x}^{i}\right)-\left(\frac{\tilde{r}}{\tilde{r_{0}}}\right)^{\frac{\mp 2}{\sqrt{D-1}}}dt^{2}+dr^{2}+\left(\frac{\tilde{r}}{\tilde{r_{0}}}\right)^{\frac{\mp 2}{\sqrt{D-1}}}\left(dx_{i}dx^{i}\right).$ (2.63)
In cosmology, such solutions exhibit scale-factor duality.
3rd kind of black hole solutions
This kind of solution in DFT involves double coordinates that are treated simultaneously, leading to a violation of the strong constraint. These solutions can be constructed using the product of self-dual solutions: $g\left(\tilde{x},x\right)=g\left(\tilde{x}\right)\times\left(g^{-1}\right)\left(x\right)$. The line element is expressed as:
$ds^{2}=g^{ij}\left(x,\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(x,\tilde{x}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(x,\tilde{x}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(x,\tilde{x}\right)dx^{i}dx^{j}$ (2.64)
Example:
Consider the following ansatz:
$dS^{2}$ $=$ $-A\left(\tilde{r},r\right)^{-1}d\tilde{t}^{2}+B\left(\tilde{r},r\right)^{-1}d\tilde{r}^{2}+C\left(\tilde{r},r\right)^{-1}\left(d\tilde{x}_{i}d\tilde{x}^{i}\right)-A\left(\tilde{r},r\right)dt^{2}+B\left(\tilde{r},r\right)dr^{2}+C\left(\tilde{r},r\right)\left(dx_{i}dx^{i}\right).$
Using this ansatz, the corresponding EOM for the graviton (LABEL:eq:_GRA_EOM) are:
$B\left[\left(\frac{\tilde{\partial}^{r}B}{B}-\frac{\tilde{\partial}^{r}A}{A}\right)\frac{\tilde{\partial}^{r}A}{A}+\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}A}{A}-2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}A}{A}\right]-\frac{1}{B}\left[\left(\frac{\partial_{r}B}{B}+\frac{\partial_{r}A}{A}\right)\frac{\partial_{r}A}{A}-\frac{\partial_{r}\partial_{r}A}{A}+2\partial_{r}d\frac{\partial_{r}A}{A}\right]$ $=$ $0,$
$B\left[\frac{1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}-\frac{1}{2}\left(\frac{\tilde{\partial}^{r}B}{B}\right)^{2}+\frac{D-2}{2}\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}B}{B}-2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}B}{B}-4\tilde{\partial}^{r}\tilde{\partial}^{r}d\right]$
$-\frac{1}{B}\left[\frac{1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}+\frac{3}{2}\left(\frac{\partial_{r}B}{B}\right)^{2}+\frac{D-2}{2}\left(\frac{\partial_{r}C}{C}\right)^{2}-\frac{\partial_{r}\partial_{r}B}{B}+2\partial_{r}d\frac{\partial_{r}B}{B}-4\partial_{r}\partial_{r}d\right]$ $=$ $0,$
$B\left[-\frac{\tilde{\partial}^{r}B}{B}\frac{\tilde{\partial}^{r}C}{C}-\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}C}{C}+\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+2\tilde{\partial}^{r}d\frac{\tilde{\partial}^{r}C}{C}\right]-\frac{1}{B}\left[-\frac{\partial_{r}B}{B}\frac{\partial_{r}C}{C}+\frac{\partial_{r}\partial_{r}C}{C}-\left(\frac{\partial_{r}C}{C}\right)^{2}-2\partial_{r}d\frac{\partial_{r}C}{C}\right]$ $=$ $0,$
Similarily, we derive the EOM for the dilaton (2.46):
$B\left(-\frac{1}{2}\left(\frac{\tilde{\partial}^{r}A}{A}\right)^{2}+\frac{1}{2}\left(\frac{\tilde{\partial}^{r}B}{B}\right)^{2}-\frac{D-2}{2}\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}-2\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}B}{B}+8\frac{\tilde{\partial}^{r}B}{B}\tilde{\partial}^{r}d+8\tilde{\partial}^{r}\tilde{\partial}^{r}d-8\tilde{\partial}^{r}d\tilde{\partial}^{r}d\right)$
$+\frac{1}{B}\left(-\frac{1}{2}\left(\frac{\partial_{r}A}{A}\right)^{2}-\frac{7}{2}\left(\frac{\partial_{r}B}{B}\right)^{2}-\frac{D-2}{2}\left(\frac{\partial_{r}C}{C}\right)^{2}+2\frac{\partial_{r}\partial_{r}B}{B}-8\frac{\partial_{r}B}{B}\partial_{r}d+8\partial_{r}\partial_{r}d-8\partial_{r}d\partial_{r}d\right)$ $=$ $0$
The solutions are expressed as:
$dS^{2}$ $=$ $-A\left(\tilde{r},r\right)^{\mp 1}d\tilde{t}^{2}+B\left(\tilde{r},r\right)^{\mp 1}d\tilde{r}^{2}+C\left(\tilde{r},r\right)^{\mp 1}\left(d\tilde{x}_{i}d\tilde{x}^{i}\right)$ (2.68)
$-A\left(\tilde{r},r\right)^{\pm}dt^{2}+B\left(\tilde{r},r\right)^{\pm}dr^{2}+C\left(\tilde{r},r\right)^{\pm}\left(dx_{i}dx^{i}\right),$
where
$A\left(\tilde{r},r\right)$ $=$ $\left(r+a\right)^{\frac{1}{\sqrt{3}}}\left(\tilde{r}+a\right)^{-\frac{1}{\sqrt{3}}},$
$B\left(\tilde{r},r\right)$ $=$ $\left(r+a\right)^{-1}\left(\tilde{r}+a\right),$
$C\left(\tilde{r},r\right)$ $=$ $\left(r+a\right)^{\sqrt{\frac{2}{3\left(D-2\right)}}}\left(\tilde{r}+a\right)^{-\sqrt{\frac{2}{3\left(D-2\right)}}},$
$d\left(\tilde{r},\tilde{\theta},r,\theta\right)$ $=$ $const.$ (2.69)
3 Space/time split double field theory
In the standard DFT, all spacetime coordinates are doubled from $D$ to $2D$, resulting in the presence of double time coordinates. This feature complicates the physical interpretation of DFT solutions. A natural resolution to this issue is to undouble the time coordinate, meaning that all fields in DFT depend on a single time coordinate while the spatial coordinates are doubled. The action for this $1+2\left(D-1\right)$-dimensional spacetime can be derived through an ADM-like decomposition of the generalized metric and the dilaton [20, 21, 22]. To perform this decomposition, we introduce the following notation: the double coordinates $X^{M}$ can be rewritten as
$X^{M}=\left(\tilde{t},t,X^{\hat{M}}\right),$ (3.70)
where the hat indices ’$\hat{M}$’ denote the spatial part of the $2D$ manifold and the tilde indices ’$\tilde{t}$’ represent the dual coordinates. Similarly, the generalized metric $\mathcal{H}^{MN}$ and the invariant metric $\eta^{MN}$ can be decomposed as
$\mathcal{H}^{MN}=\left(\begin{array}[]{ccc}\mathcal{H}^{\tilde{t}\tilde{t}}&0&0\\ 0&\mathcal{H}^{tt}&0\\ 0&0&\mathcal{H}^{\hat{M}\hat{N}}\end{array}\right),\qquad\eta^{MN}=\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&\eta^{\hat{M}\hat{N}}\end{array}\right),$ (3.71)
where we set the generalized shift vector to zero, thus $\mathcal{H}^{tM}$=$\mathcal{H}^{\tilde{t}M}$=0, $\mathcal{H}^{tt}=-n^{-2}$, $\mathcal{H}^{\tilde{t}\tilde{t}}=-n^{2}$ and $n$ is the lapse function. The dilaton can then be rewritten as
$e^{-2d}=ne^{-2\hat{d}}.$ (3.72)
Substituting the decompositions (3.71) and (3.72) back into the DFT action, we obtain:
$S_{\mathrm{DFT}}=\int d^{D}xd^{D}\tilde{x}\mathcal{L}_{\mathrm{DFT}}\left(d,\mathcal{H}_{MN}\right),$ (3.73)
where
$\mathcal{L}_{\mathrm{DFT}}\left(d,\mathcal{H}_{MN}\right)$ $=$ $e^{-2d}\left(\frac{1}{8}\mathcal{H}^{MN}\partial_{M}\mathcal{H}^{KL}\partial_{N}\mathcal{H}_{KL}-\frac{1}{2}\mathcal{H}^{MN}\partial_{N}\mathcal{H}^{KL}\partial_{L}\mathcal{H}_{MK}\right.$ (3.74)
$\left.-2\partial_{M}d\partial_{N}\mathcal{H}^{MN}+4\mathcal{H}^{MN}\partial_{M}d\partial_{N}d\right).$
By removing the dual time-like coordinate through the constraint $\frac{\partial}{\partial\tilde{t}}\left(\cdots\right)=0$, we obtain:
[…]
where
[…]
This implies that we replace $d$ as $\hat{d}-\frac{1}{2}\log n$ and $\mathcal{H}_{MN}$ with $\mathcal{H}_{\hat{M}\hat{N}}$ in the Lagrangian (3.74). This action is consistent with [21]. After a brief calculation, it also shows that
$\mathcal{L}_{\mathrm{DFT}}\left(\hat{d}-\frac{1}{2}\log n,\mathcal{H}_{\hat{M}\hat{N}}\right)-n^{-1}e^{-2d}\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}n\partial_{\hat{N}}n=ne^{-2d}\mathcal{R}\left(\hat{d},\mathcal{H}_{\hat{M}\hat{N}}\right)-b_{1},$ (3.77)
where $b_{1}$ is a total derivative term (boundary term):
$b_{1}=-\partial_{\hat{M}}\left(ne^{-2\hat{d}}\left[\partial_{\hat{N}}\mathcal{H}^{\hat{M}\hat{N}}-4\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{N}}\hat{d}\right]\right).$ (3.78)
Finally, the $1+2\left(D-1\right)$-dimensional DFT action is given by
[…]
Since all decomposed fields of DFT only depend on $\left(t,X^{\hat{M}}\right)$:
$\mathcal{H}^{\hat{M}\hat{N}}\left(t,X^{\hat{M}}\right),\qquad\hat{d}\left(t,X^{\hat{M}}\right),\qquad n\left(t,X^{\hat{M}}\right).$ (3.80)
and the spacetime action is an integral over a $1+2\left(D-1\right)$-dimensional spacetime, the dual timelike direction $\tilde{t}$ has no physical meaning. This implies that we can safely ignore the dual timelike direction $\tilde{t}$, and the action (3.79) can be effectively considered as the action for a $1+2\left(D-1\right)$ spacetime with the following ansatz:
$ds^{2}=-n^{2}dt^{2}+\mathcal{H}_{\hat{M}\hat{N}}dX^{\hat{M}}dX^{\hat{N}},$ (3.81)
which can be expanded as
$ds^{2}$ $=$ $-n^{2}dt^{2}+\mathcal{H}_{\hat{\mathbf{1}}\hat{\mathbf{1}}}dX^{\hat{\mathbf{1}}}dX^{\hat{\mathbf{1}}}+\mathcal{H}_{\hat{\mathbf{1}}\hat{\mathbf{2}}}dX^{\hat{\mathbf{1}}}dX^{\hat{\mathbf{2}}}+\mathcal{H}_{\hat{\mathbf{2}}\hat{\mathbf{1}}}dX^{\hat{\mathbf{2}}}dX^{\hat{\mathbf{1}}}+\mathcal{H}_{\hat{\mathbf{2}}\hat{\mathbf{2}}}dX^{\hat{\mathbf{2}}}dX^{\hat{\mathbf{2}}}$ (3.82)
$=$ $-n^{2}dt^{2}+g^{ij}d\tilde{x}_{i}d\tilde{x}_{j}-g^{ik}b_{kj}d\tilde{x}_{i}dx^{j}+b_{ik}g^{kj}dx^{i}d\tilde{x}_{j}+\left(g_{ij}-b_{ik}g^{kl}b_{lj}\right)dx^{i}dx^{j}.$
For simplicity in the rest of the paper, we set the Kalb-Ramond field to zero, so the line element becomes:
$ds^{2}=-n^{2}dt^{2}+\mathcal{H}_{\hat{\mathbf{1}}\hat{\mathbf{1}}}dX^{\hat{\mathbf{1}}}dX^{\hat{\mathbf{1}}}+\mathcal{H}_{\hat{\mathbf{2}}\hat{\mathbf{2}}}dX^{\hat{\mathbf{2}}}dX^{\hat{\mathbf{2}}}=-n^{2}dt^{2}+g^{ij}d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}dx^{i}dx^{j}.$ (3.83)
Thus, our aim is to calculate the black hole solutions for this ansatz in the rest of the paper.
To obtain the EOM, we can also substitute the field decompositions (3.71) and (3.72) into the equations (2.25). The dilatonic EOM (2.46) then becomes:
$\frac{1}{8}\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}\mathcal{H}^{\hat{K}\hat{L}}\partial_{\hat{N}}\mathcal{H}_{\hat{K}\hat{L}}-\frac{1}{2}\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}\mathcal{H}^{\hat{K}\hat{L}}\partial_{\hat{K}}\mathcal{H}_{\hat{N}\hat{L}}-\partial_{\hat{M}}\partial_{\hat{N}}\mathcal{H}^{\hat{M}\hat{N}}$
$-4\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}d\partial_{\hat{N}}d+4\partial_{\hat{M}}\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{N}}d+4\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}\partial_{\hat{N}}d$
$+\frac{1}{4}\frac{1}{n^{2}}\partial_{t}n^{-2}\partial_{t}n^{2}+\partial_{t}\partial_{t}n^{-2}-4\partial_{t}n^{-2}\partial_{t}d-\frac{1}{8}\frac{1}{n^{2}}\partial_{t}\mathcal{H}^{\hat{K}\hat{L}}\partial_{t}\mathcal{H}_{\hat{K}\hat{L}}+\frac{1}{4}\mathcal{H}^{\hat{M}\hat{N}}\partial_{\hat{M}}n^{-2}\partial_{\hat{N}}n^{2}$ $=$ $0$ (3.84)
The gravitational EOM (LABEL:eq:_GRA_EOM) are given by:
$\mathcal{R}_{tt}=\mathcal{K}_{tt}-S_{\quad t}^{P}\mathcal{K}_{PQ}S_{\quad t}^{Q}$ $=$ $0,$
$\mathcal{R}_{\hat{M}\hat{N}}=\mathcal{K}_{\hat{M}\hat{N}}-S_{\quad\hat{M}}^{P}\mathcal{K}_{PQ}S_{\quad\hat{N}}^{Q}$ $=$ $0$ (3.85)
For simplicity, we will continue to use $d$ instead of $\hat{d}$ in the rest of the paper.
3.1 Black hole solutions
To calculate the black hole solutions of space/time split DFT, we utilize the following ansatz:
$ds^{2}=-n^{2}dt^{2}+\mathcal{H}_{\hat{M}\hat{N}}dX^{\hat{M}}dX^{\hat{N}},$ (3.86)
where we assume the generalized metric with the vanishing Kalb-Ramond field takes the form:
$\mathcal{H}_{\hat{M}\left(i\right)\hat{N}\left(j\right)}\equiv\left(\begin{array}[]{cc}\mathcal{H}_{\hat{\mathbf{1}}\left(i\right)\hat{\mathbf{1}}\left(j\right)}&0\\ 0&\mathcal{H}_{\hat{\mathbf{2}}\left(i\right)\hat{\mathbf{2}}\left(j\right)}\end{array}\right)=\left(\begin{array}[]{cc}g^{ij}&0\\ 0&g_{ij}\end{array}\right),$ (3.87)
where $g_{ij}$ denotes the spatial section of spacetime. Substituting this ansatz into the gravitational EOM (3.85), the non-vanishing generalized Ricci scalar can be given by $\mathcal{R}_{\hat{M}\left(i\right)\hat{N}\left(j\right)}$ and $\mathcal{R}_{tt}$ which are given as follows:
$\mathcal{R}_{\hat{M}\left(i\right)\hat{N}\left(j\right)}=\mathcal{K}_{\hat{M}\hat{N}}-S_{\quad\hat{M}}^{P}\mathcal{K}_{PQ}S_{\quad\hat{N}}^{Q}\equiv\left(\begin{array}[]{cc}\mathcal{R}_{\hat{\mathbf{1}}\left(i\right)\hat{\mathbf{1}}\left(j\right)}&\mathcal{R}_{\hat{\mathbf{1}}\left(i\right)\hat{\mathbf{2}}\left(j\right)}\\ \mathcal{R}_{\hat{\mathbf{2}}\left(i\right)\hat{\mathbf{1}}\left(j\right)}&\mathcal{R}_{\hat{\mathbf{2}}\left(i\right)\hat{\mathbf{2}}\left(j\right)}\end{array}\right),$ (3.88)
and
$\mathcal{R}_{tt}$ $=$ $\frac{1}{4}\partial_{t}g_{ij}\partial_{t}g^{ij}-\frac{1}{4}\partial_{t}n^{2}\partial_{t}n^{-2}+\frac{1}{4}\partial_{t}\left(n^{-2}\partial_{t}n^{2}\right)$ (3.89)
$-\frac{1}{4}n^{4}\partial_{t}\left(n^{-2}\partial_{t}n^{-2}\right)-\frac{1}{2}\partial_{t}d\left(n^{-2}\partial_{t}n^{2}\right)+2\partial_{t}\partial_{t}d$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}n^{2}\right)-\frac{1}{4}n^{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}n^{-2}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}n^{2}\right)+\frac{1}{2}n^{4}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}n^{-2}\right)$
$+\frac{1}{4}\partial_{i}\left(g^{ij}\partial_{j}n^{2}\right)-\frac{1}{4}n^{4}\partial_{i}\left(g^{ij}\partial_{j}n^{-2}\right)-\frac{1}{2}\partial_{i}d\left(g^{ij}\partial_{j}n^{2}\right)+\frac{1}{2}n^{4}\partial_{i}d\left(g^{ij}\partial_{j}n^{-2}\right),$
Then, $\mathcal{R}_{\hat{M}\left(i\right)\hat{N}\left(j\right)}$ can be expanded as:
$\mathcal{R}_{\hat{\mathbf{1}}\left(p\right)\hat{\mathbf{1}}\left(q\right)}$ $=$ $\frac{1}{4}\tilde{\partial}^{p}n^{-2}\tilde{\partial}^{q}n^{2}-\frac{1}{4}g^{ip}\partial_{i}n^{2}\partial_{j}n^{-2}g^{jq}+\frac{1}{4}\partial_{t}\left(n^{-2}\partial_{t}g^{pq}\right)+\frac{1}{4}g^{ip}\partial_{t}\left(n^{-2}\partial_{t}n^{2}\right)g^{jq}$ (3.90)
$-\frac{1}{2}\partial_{t}d\left(n^{-2}\partial_{t}g^{pq}\right)+\frac{1}{2}g^{ip}\partial_{t}d\left(n^{-2}\partial_{t}g_{ij}\right)g^{jq}$
$+\frac{1}{4}\tilde{\partial}^{p}g^{ij}\tilde{\partial}^{q}g_{ij}-\frac{1}{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}g^{pq}\right)-\frac{1}{4}\partial_{i}\left(g^{ij}\partial_{j}g^{pq}\right)$
$-\frac{1}{4}\tilde{\partial}^{q}g_{ij}\tilde{\partial}^{j}g^{pi}-\frac{1}{4}\tilde{\partial}^{p}g_{ij}\tilde{\partial}^{j}g^{qi}+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\tilde{\partial}^{q}g^{pj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\tilde{\partial}^{p}g^{qj}\right)$
$+\frac{1}{4}\partial_{i}\left(g^{jp}\partial_{j}g^{iq}\right)+\frac{1}{4}\partial_{i}\left(g^{jq}\partial_{j}g^{ip}\right)-\frac{1}{8}g^{ip}\partial_{i}g_{kl}\partial_{j}g^{kl}g^{jq}-\frac{1}{8}g^{ip}\partial_{i}g^{kl}\partial_{j}g_{kl}g^{jq}$
$+\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{lk}\tilde{\partial}^{k}g_{ij}\right)g^{jq}+\frac{1}{4}g^{ip}\partial_{l}\left(g^{lk}\partial_{k}g_{ij}\right)g^{jq}+\frac{1}{4}g^{ip}\partial_{j}g^{kl}\partial_{l}g_{ik}g^{jq}+\frac{1}{4}g^{ip}\partial_{i}g^{kl}\partial_{l}g_{jk}g^{jq}$
$-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\partial_{j}g_{ik}\right)g^{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\partial_{i}g_{jk}\right)g^{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{ki}\tilde{\partial}^{k}g_{lj}\right)g^{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{kj}\tilde{\partial}^{k}g_{li}\right)g^{jq}$
$+\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}g^{pq}\right)+\frac{1}{2}\partial_{i}d\left(g^{ij}\partial_{j}g^{pq}\right)+2\tilde{\partial}^{p}\tilde{\partial}^{q}d$
$-\frac{1}{2}\partial_{i}d\left(g^{jp}\partial_{j}g^{iq}\right)-\frac{1}{2}\partial_{i}d\left(g^{jq}\partial_{j}g^{ip}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ji}\tilde{\partial}^{q}g^{pj}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ji}\tilde{\partial}^{p}g^{qj}\right)$
$-\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{lk}\tilde{\partial}^{k}g_{ij}\right)g^{jq}-\frac{1}{2}g^{ip}\partial_{l}d\left(g^{lk}\partial_{k}g_{ij}\right)g^{jq}-2g^{ip}\partial_{i}\partial_{j}dg^{jq}$
$+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\partial_{j}g_{ik}\right)g^{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\partial_{i}g_{jk}\right)g^{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{ki}\tilde{\partial}^{k}g_{lj}\right)g^{jq}+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{kj}\tilde{\partial}^{k}g_{li}\right)g^{jq},$
$\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{2}}\left(q\right)}$ $=$ $\frac{1}{4}\partial_{p}n^{2}\partial_{q}n^{-2}-\frac{1}{4}g_{pp}\tilde{\partial}^{p}n^{2}\tilde{\partial}^{q}n^{-2}g_{qq}+\frac{1}{4}\partial_{t}\left(n^{-2}\partial_{t}g_{pq}\right)+\frac{1}{4}g_{ip}\partial_{t}\left(n^{-2}\partial_{t}g^{ij}\right)g_{jq}$ (3.91)
$-\frac{1}{2}g_{ip}\partial_{t}d\left(n^{-2}\partial_{t}g^{ij}\right)g_{jq}+\frac{1}{2}\partial_{t}d\left(n^{-2}\partial_{t}g_{pq}\right)$
$+\frac{1}{4}\partial_{p}g^{ij}\partial_{q}g_{ij}-\frac{1}{4}\tilde{\partial}^{i}\left(g_{ij}\tilde{\partial}^{j}g_{pq}\right)-\frac{1}{4}\partial_{i}\left(g^{ij}\partial_{j}g_{pq}\right)$
$-\frac{1}{4}\partial_{q}g^{ij}\partial_{j}g_{pi}-\frac{1}{4}\partial_{p}g^{ij}\partial_{j}g_{qi}+\frac{1}{4}\partial_{i}\left(g^{ji}\partial_{q}g_{pj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\partial_{p}g_{qj}\right)$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{jp}\tilde{\partial}^{j}g_{iq}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g_{jq}\tilde{\partial}^{j}g_{ip}\right)-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{j}g^{kl}g_{jq}-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g^{kl}\tilde{\partial}^{j}g_{kl}g_{jq}$
$+\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{lk}\tilde{\partial}^{k}g^{ij}\right)g_{jq}+\frac{1}{4}g_{ip}\partial_{l}\left(g^{lk}\partial_{k}g^{ij}\right)g_{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{j}g_{kl}\tilde{\partial}^{l}g^{ik}g_{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{l}g^{jk}g_{jq}$
$-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\tilde{\partial}^{j}g^{ik}\right)g_{jq}-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\tilde{\partial}^{i}g^{jk}\right)g_{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{ki}\partial_{k}g^{lj}\right)g_{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{kj}\partial_{k}g^{li}\right)g_{jq}$
$+\frac{1}{2}\tilde{\partial}^{i}d\left(g_{ij}\tilde{\partial}^{j}g_{pq}\right)+\frac{1}{2}\partial_{i}d\left(g^{ij}\partial_{j}g_{pq}\right)+2\partial_{p}\partial_{q}d$
$-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{jp}\tilde{\partial}^{j}g_{iq}\right)-\frac{1}{2}\tilde{\partial}^{i}d\left(g_{jq}\tilde{\partial}^{j}g_{ip}\right)-\frac{1}{2}\partial_{i}d\left(g^{ji}\partial_{q}g_{pj}\right)-\frac{1}{2}\partial_{i}d\left(g^{ji}\partial_{p}g_{qj}\right)$
$-\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{lk}\tilde{\partial}^{k}\tilde{g}^{ij}\right)g_{jq}-\frac{1}{2}g_{ip}\partial_{l}d\left(g^{lk}\partial_{k}g^{ij}\right)g_{jq}-2g_{ip}\tilde{\partial}^{i}\tilde{\partial}^{j}dg_{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\tilde{\partial}^{j}g^{ik}\right)g_{jq}+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\tilde{\partial}^{i}g^{jk}\right)g_{jq}$
$+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{ki}\partial_{k}g^{lj}\right)g_{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{kj}\partial_{k}g^{li}\right)g_{jq},$
$\mathcal{R}_{\hat{\mathbf{1}}\left(p\right)\hat{\mathbf{2}}\left(q\right)}$ $=$ $\frac{1}{8}\tilde{\partial}^{p}n^{2}\partial_{q}n^{-2}+\frac{1}{8}\tilde{\partial}^{p}n^{-2}\partial_{q}n^{2}$ (3.92)
$+\frac{1}{8}\tilde{\partial}^{p}g_{ij}\partial_{q}g^{ij}+\frac{1}{8}\tilde{\partial}^{p}g^{ij}\partial_{q}g_{ij}-\frac{1}{4}\partial_{q}g_{ij}\tilde{\partial}^{j}g^{pi}-\frac{1}{4}\tilde{\partial}^{p}g^{ij}\partial_{j}g_{qi}$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\partial_{q}g^{pj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\tilde{\partial}^{p}g_{qj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g^{jp}\partial_{j}g_{iq}\right)+\frac{1}{4}\partial_{i}\left(g_{jq}\tilde{\partial}^{j}g^{ip}\right)$
$-\frac{1}{8}g^{ip}\partial_{i}g_{kl}\tilde{\partial}^{j}g^{kl}g_{jq}-\frac{1}{8}g^{ip}\partial_{i}g^{kl}\tilde{\partial}^{j}g_{kl}g_{jq}+\frac{1}{4}g^{ip}\partial_{i}g_{kl}\tilde{\partial}^{l}g^{jk}g_{jq}+\frac{1}{4}g^{ip}\tilde{\partial}^{j}g^{kl}\partial_{l}g_{ik}g_{jq}$
$-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g_{kl}\partial_{i}g^{jk}\right)g_{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g^{kl}\tilde{\partial}^{j}g_{ik}\right)g_{jq}-\frac{1}{4}g^{ip}\partial_{l}\left(g_{ki}\tilde{\partial}^{k}g^{lj}\right)g_{jq}-\frac{1}{4}g^{ip}\tilde{\partial}^{l}\left(g^{kj}\partial_{k}g_{li}\right)g_{jq}$
$+2\tilde{\partial}^{p}\partial_{q}d-\tilde{\partial}^{i}d\left(g^{jp}\partial_{j}g_{iq}\right)-\partial_{i}d\left(g_{jq}\tilde{\partial}^{j}g^{ip}\right)$
~$-\tilde{\partial}^{i}d\left(g_{ji}\partial_{q}g^{pj}\right)-\partial_{i}d\left(g^{ji}\tilde{\partial}^{p}g_{qj}\right)-2g^{ip}\partial_{i}\tilde{\partial}^{j}dg_{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g_{kl}\partial_{i}g^{jk}\right)g_{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g^{kl}\tilde{\partial}^{j}g_{ik}\right)g_{jq}$
$+\frac{1}{2}g^{ip}\tilde{\partial}^{l}d\left(g^{kj}\partial_{k}g_{li}\right)g_{jq}+\frac{1}{2}g^{ip}\partial_{l}d\left(g_{ki}\tilde{\partial}^{k}g^{lj}\right)g_{jq},$
$\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{1}}\left(q\right)}$ $=$ $\frac{1}{8}\partial_{p}n^{2}\tilde{\partial}^{q}n^{-2}+\frac{1}{8}\partial_{p}n^{-2}\tilde{\partial}^{q}n^{2}$ (3.93)
$+\frac{1}{8}\partial_{p}g_{ij}\tilde{\partial}^{q}g^{ij}+\frac{1}{8}\partial_{p}g^{ij}\tilde{\partial}^{q}g_{ij}-\frac{1}{4}\partial_{p}g_{ij}\tilde{\partial}^{j}g^{qi}-\frac{1}{4}\tilde{\partial}^{q}g^{ij}\partial_{j}g_{pi}$
$+\frac{1}{4}\tilde{\partial}^{i}\left(g_{ji}\partial_{p}g^{qj}\right)+\frac{1}{4}\partial_{i}\left(g^{ji}\tilde{\partial}^{q}g_{pj}\right)+\frac{1}{4}\tilde{\partial}^{i}\left(g^{jq}\partial_{j}g_{ip}\right)+\frac{1}{4}\partial_{i}\left(g_{jp}\tilde{\partial}^{j}g^{iq}\right)$
$-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g_{kl}\partial_{j}g^{kl}g^{jq}-\frac{1}{8}g_{ip}\tilde{\partial}^{i}g^{kl}\partial_{j}g_{kl}g^{jq}+\frac{1}{4}g_{ip}\partial_{j}g_{kl}\tilde{\partial}^{l}g^{ik}g^{jq}+\frac{1}{4}g_{ip}\tilde{\partial}^{i}g^{kl}\partial_{l}g_{jk}g^{jq}$
$-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g_{kl}\partial_{j}g^{ik}\right)g^{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g^{kl}\tilde{\partial}^{i}g_{jk}\right)g^{jq}-\frac{1}{4}g_{ip}\tilde{\partial}^{l}\left(g^{ki}\partial_{k}g_{lj}\right)g^{jq}-\frac{1}{4}g_{ip}\partial_{l}\left(g_{kj}\tilde{\partial}^{k}g^{li}\right)g^{jq}$
$+2\partial_{p}\tilde{\partial}^{q}d-\tilde{\partial}^{i}d\left(g^{jq}\partial_{j}g_{ip}\right)-\partial_{i}d\left(g_{jp}\tilde{\partial}^{j}g^{iq}\right)$
~$-\tilde{\partial}^{i}d\left(g_{ji}\partial_{p}g^{qj}\right)-\partial_{i}d\left(g^{ji}\tilde{\partial}^{q}g_{pj}\right)-2g_{ip}\tilde{\partial}^{i}\partial_{j}dg^{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g_{kl}\partial_{j}g^{ik}\right)g^{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g^{kl}\tilde{\partial}^{i}g_{jk}\right)g^{jq}$
$+\frac{1}{2}g_{ip}\tilde{\partial}^{l}d\left(g^{ki}\partial_{k}g_{lj}\right)g^{jq}+\frac{1}{2}g_{ip}\partial_{l}d\left(g_{kj}\tilde{\partial}^{k}g^{li}\right)g^{jq}.$
On the other hand, the EOM for the dilaton (3.84) is expressed as:
$\frac{1}{8}g_{ij}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{j}g^{kl}+\frac{1}{8}g_{ij}\tilde{\partial}^{i}g^{kl}\tilde{\partial}^{j}g_{kl}+\frac{1}{8}g^{ij}\partial_{i}g_{kl}\partial_{j}g^{kl}+\frac{1}{8}g^{ij}\partial_{i}g^{kl}\partial_{j}g_{kl}$
$-\frac{1}{2}g_{ij}\tilde{\partial}^{i}g_{kl}\tilde{\partial}^{k}g^{jl}-\frac{1}{2}g^{ij}\partial_{i}g^{kl}\partial_{k}g_{jl}-\tilde{\partial}^{i}\tilde{\partial}^{j}g_{ij}-\partial_{i}\partial_{j}g^{ij}-4g_{ij}\tilde{\partial}^{i}d\tilde{\partial}^{j}d-4g^{ij}\partial_{i}d\partial_{j}d$
$+4\tilde{\partial}^{i}g_{ij}\tilde{\partial}^{j}d+4\partial_{i}g^{ij}\partial_{j}d+4g_{ij}\tilde{\partial}^{i}\tilde{\partial}^{j}d+4g^{ij}\partial_{i}\partial_{j}d+\frac{1}{4}g_{ij}\tilde{\partial}^{i}n^{2}\tilde{\partial}^{j}n^{-2}+\frac{1}{4}g^{ij}\partial_{i}n^{2}\partial_{j}n^{-2}$
$-\frac{1}{4}n^{-2}\partial_{t}g_{kl}\partial_{t}g^{kl}+\frac{1}{4}n^{-2}\partial_{t}n^{2}\partial_{t}n^{-2}+\partial_{t}\partial_{t}n^{-2}-4\partial_{t}n^{-2}\partial_{t}d=0.$ (3.94)
Next, we consider the $D=1+\left(2+2\right)$ dimensional black hole line element:
$dS^{2}=-n^{2}dt^{2}+\mathcal{H}_{\hat{M}\hat{N}}dX^{\hat{M}}dX^{\hat{N}}=-n\left(r\right)^{2}dt^{2}+g^{ij}d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}dx^{i}dx^{j},$ (3.95)
where the strong constraint $\eta_{\hat{M}\hat{N}}\partial^{\hat{M}}\partial^{\hat{N}}\left(\cdots\right)=0$ is imposed, and the metric assumes the form:
$g_{ij}\left(r\right)=\left(\begin{array}[]{cc}B\left(r\right)&0\\ 0&C\left(r\right)\end{array}\right).$ (3.96)
Furthermore, due to symmetries among generalized Ricci scalars $\mathcal{R}_{M\left(p\right)N\left(q\right)}$,
$\mathcal{R}_{\hat{\mathbf{1}}\left(p\right)\mathbf{1}\left(q\right)}$ $\underleftrightarrow{g^{\bullet\bullet}\leftrightarrow g_{\bullet\bullet},\quad\tilde{\partial}^{\bullet}\leftrightarrow\partial_{\bullet}}$ $\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{2}}\left(q\right)},$
$\mathcal{R}_{\hat{\mathbf{1}}\left(p\right)\hat{\mathbf{2}}\left(q\right)}$ $\underleftrightarrow{g^{\bullet\bullet}\leftrightarrow g_{\bullet\bullet},\quad\tilde{\partial}^{\bullet}\leftrightarrow\partial_{\bullet}}$ $\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{1}}\left(q\right)},$ (3.97)
we only need to calculate one set of them. Here, we compute $\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{2}}\left(q\right)}$ and $\mathcal{R}_{\hat{\mathbf{2}}\left(p\right)\hat{\mathbf{1}}\left(q\right)}$. Consequently, the gravitational EOM (3.85) become:
$\mathcal{R}_{tt}$ $=$ $nB\left[\left(\frac{\tilde{\partial}^{r}B}{B}-\frac{\tilde{\partial}^{r}n}{n}\right)\tilde{\partial}^{r}n+\tilde{\partial}^{r}\tilde{\partial}^{r}n-2\tilde{\partial}^{r}d\tilde{\partial}^{r}n\right]$
$-n\frac{1}{B}\left[\left(\frac{\partial_{r}B}{B}+\frac{\partial_{r}n}{n}\right)\partial_{r}n-\partial_{r}\partial_{r}n+2\partial_{r}d\partial_{r}n\right]=0,$
$\mathcal{R}_{\hat{\mathbf{2}}\left(r\right)\hat{\mathbf{2}}\left(r\right)}$ $=$ $B\left[B\left(\frac{\tilde{\partial}^{r}n}{n}\right)^{2}+\frac{1}{4}B\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}+\frac{1}{2}\tilde{\partial}^{r}\tilde{\partial}^{r}B-\frac{1}{4}\frac{\left(\tilde{\partial}^{r}B\right)^{2}}{B}-\tilde{\partial}^{r}d\tilde{\partial}^{r}B-2B\tilde{\partial}^{r}\tilde{\partial}^{r}d\right]$
$-\frac{1}{B}\left[B\left(\frac{\partial_{r}n}{n}\right)^{2}+\frac{1}{4}B\left(\frac{\partial_{r}C}{C}\right)^{2}-\frac{1}{2}\partial_{r}\partial_{r}B+\frac{3}{4}\frac{\left(\partial_{r}B\right)^{2}}{B}+\partial_{r}d\partial_{r}B-2B\partial_{r}\partial_{r}d\right]=0,$
$\mathcal{R}_{\hat{\mathbf{2}}\left(\varphi\right)\hat{\mathbf{2}}\left(\varphi\right)}$ $=$ $\frac{1}{2}B\left[-\frac{\tilde{\partial}^{r}B}{B}\tilde{\partial}^{r}C-\tilde{\partial}^{r}\tilde{\partial}^{r}C+\frac{\left(\tilde{\partial}^{r}C\right)^{2}}{C}+2\tilde{\partial}^{r}d\tilde{\partial}^{r}C\right]$ (3.98)
$-\frac{1}{2}\frac{1}{B}\left[-\frac{\partial_{r}B}{B}\partial_{r}C+\partial_{r}\partial_{r}C-\frac{\left(\partial_{r}C\right)^{2}}{C}-2\partial_{r}d\partial_{r}C\right]=0.$
Moreover, to compute the black hole solution in non-critical string theory, we must incorporate the cosmological constant into the action. This constant does not break the $O\left(d,d\right)$ symmetry and the diffeomorphism of the action:
[…]
Therefore, the EOM for the dilaton (3.84) is expressed as:
$B\left[-\left(\frac{\tilde{\partial}^{r}n}{n}\right)^{2}+\frac{1}{4}\left(\frac{\tilde{\partial}^{r}B}{B}\right)^{2}-\frac{1}{4}\left(\frac{\tilde{\partial}^{r}C}{C}\right)^{2}-\frac{\tilde{\partial}^{r}\tilde{\partial}^{r}B}{B}+4\frac{\tilde{\partial}^{r}B}{B}\tilde{\partial}^{r}d-4\tilde{\partial}^{r}d\tilde{\partial}^{r}d+4\tilde{\partial}^{r}\tilde{\partial}^{r}d\right]$
$+\frac{1}{B}\left[-\left(\frac{\partial_{r}n}{n}\right)^{2}-\frac{7}{4}\left(\frac{\partial_{r}B}{B}\right)^{2}-\frac{1}{4}\left(\frac{\partial_{r}C}{C}\right)^{2}+\frac{\partial_{r}\partial_{r}B}{B}-4\frac{\partial_{r}B}{B}\partial_{r}d-4\partial_{r}d\partial_{r}d+4\partial_{r}\partial_{r}d\right]$ $=$ $-\lambda^{2}.$
At first, we attempt to derive a solution corresponding to the 3rd kind of solution found in standard DFT as discussed in the previous section. The solution, which violates constraints, is given by:
$ds^{2}=\left[-n\left(r,\tilde{r}\right)^{2}dt^{2}+dr^{2}+dx^{2}\right]+\left[d\tilde{r}^{2}+d\tilde{x}^{2}\right],$ (3.101)
where
$n\left(r,\tilde{r}\right)$ $=$ $\left(\tanh\left(\frac{\sqrt{2}\lambda}{4}\tilde{r}\right)\right)^{\mp 1}\left(\tanh\left(\frac{\sqrt{2}\lambda}{4}r\right)\right)^{\pm 1},$
$d\left(r,\tilde{r}\right)$ $=$ $-\frac{1}{2}\log\left(\sinh\left(\frac{\sqrt{2}\lambda}{2}\tilde{r}\right)\sinh\left(\frac{\sqrt{2}\lambda}{2}r\right)\right).$ (3.102)
This configuration represents a double black string solution. However, it violates the strong constraint due to its dependence on dual coordinates $\left(r,\tilde{r}\right)$. Further investigation into this solution is postponed until a compelling reason to relax the strong constraint and interpret dual coordinates emerges. Alternatively, we impose the strong constraint and derive two types of solutions for the EOM (3.98) and (LABEL:eq:dilaton_eom), corresponding to the 1st kind of solution discussed in the previous section. The solution is
$ds^{2}=\left[-n\left(r\right)^{2}dt^{2}+B\left(r\right)dr^{2}+C\left(r\right)d\varphi^{2}\right]+\left[B\left(r\right)^{-1}d\tilde{r}^{2}+C\left(r\right)^{-1}d\tilde{\varphi}^{2}\right],$ (3.103)
where
$n\left(r\right)=\left(\sqrt{\left(1-\frac{M}{r}\right)^{a}}\right)^{\pm 1},\qquad B\left(r\right)=\left(1-\frac{M}{r}\right)^{-c}\frac{1}{\lambda^{2}r^{2}},\qquad C\left(r\right)=\left(\left(1-\frac{M}{r}\right)^{b}\frac{1}{\lambda^{2}}\right)^{\pm 1},$
$d=\mathrm{const},\qquad a^{2}+b^{2}=1,\qquad c=1.$ (3.104)
The choices of $\pm$ for $n\left(r\right)$ and $C\left(r\right)$ do not need to be the same simultaneously. The dual solution is given by:
$ds^{2}=\left[-n\left(\tilde{r}\right)^{2}dt^{2}+B\left(\tilde{r}\right)^{-1}d\tilde{r}^{2}+C\left(\tilde{r}\right)^{-1}d\tilde{\varphi}^{2}\right]+\left[B\left(\tilde{r}\right)dr^{2}+C\left(\tilde{r}\right)d\varphi^{2}\right]$ (3.105)
where
$n\left(\tilde{r}\right)=\left(\sqrt{\left(1-\frac{M}{\tilde{r}}\right)^{a}}\right)^{\mp 1},\qquad B\left(\tilde{r}\right)=\left(1-\frac{M}{\tilde{r}}\right)^{c}\lambda^{2}\tilde{r}^{2},\qquad C\left(\tilde{r}\right)=\left(\left(1-\frac{M}{\tilde{r}}\right)^{-b}\lambda^{2}\right)^{\mp 1},$
$d=\mathrm{const},\qquad a^{2}+b^{2}=1,\qquad c=1.$ (3.106)
A similar solution in the low-energy effective action was studied in reference [45]. Note that our solutions apply to both compact and noncompact $\varphi$.
For simplicity, we analyze only one of these solutions, specifically:
$ds^{2}$ $=$ $\left[-\left(1-\frac{M}{r}\right)^{a}dt^{2}+\left(1-\frac{M}{r}\right)^{-c}\frac{1}{\lambda^{2}r^{2}}dr^{2}+\left(1-\frac{M}{r}\right)^{b}\frac{1}{\lambda^{2}}d\varphi^{2}\right]$ (3.107)
$+\left[\left(1-\frac{M}{r}\right)^{c}\lambda^{2}r^{2}d\tilde{r}^{2}+\left(1-\frac{M}{r}\right)^{-b}\lambda^{2}d\tilde{\varphi}^{2}\right].$
To determine its physical interpretation, we introduce two key results:
A. Generalized scalar curvature
To investigate the singularities of this metric, we need to consider the scalar curvature within the framework of this theory. The geometry of DFT is structured around generalized diffeomorphisms. Therefore, the connection must incorporate covariant derivatives that respect these symmetries. This connection can be defined analogously to the Levi-Civita connection in Riemannian geometry, ensuring compatibility with the generalized metric and the $O\left(d,d\right)$ structure. Using this connection, the generalized Ricci curvature is derived under the condition of vanishing torsion. However, the generalized Ricci scalar $\mathcal{R}$ and tensor $\mathcal{R}_{MN}$ are constrained to vanish due to satisfying the equations of motion for the graviton and dilaton: $\mathcal{R}=0$ and $\mathcal{R}_{MN}=0$. This property complicates the measurement of curvature, even in spacetime singularities within DFT [41]. In our solutions, incorporating the cosmological constant in DFT leads to the generalized Ricci scalar being
$\mathcal{R}=-\lambda^{2},$ (3.108)
indicating that our solutions represent AdS vacuum solutions in double coordinates. Importantly, DFT possesses AdS vacuum solutions due to the generalized Ricci scalar includes dilatonic effects.
B. Reduction to low-energy effective action
To impose the strong constraint ~$\tilde{\partial}\left(\ldots\right)=0$, requiring all fields to be independent of the dual coordinates, the DFT action simplifies to the traditional low-energy effective action:
[…]
Alternatively, this also can be expressed as:
|~$\left.\mathcal{R}\right|_{\tilde{\partial}=0}=\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}-\frac{1}{12}H_{\mu\nu\rho}H^{\mu\nu\rho}\right].$ (3.110)
While integrating out the dual coordinates $\tilde{x}$ only yields an overall constant in the action, it may seem negligible. However, to retain the novel insights of DFT, we preserve this term and interpret the generalized metric as the complete metric of spacetime. Moreover, this reduction implies that the undoubled geometry described by $x$ can be entirely governed by the traditional low-energy effective action, without the need for any additional theories. This indicates that the traditional interpretation of black holes in the low-energy effective action also applies to DFT black holes. Thus, the Kretschmann scalar for the undoubled spacetime (spanned by $x$) of the generalized metric (3.107) is given by
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}\left(1-\frac{M}{r}\right)^{2c}\left(a^{4}M^{2}+2a^{3}M\left(cM-2r\right)+a^{2}\left(b^{2}M^{2}+\left(cM-2r\right)^{2}\right)+b^{2}\left(bM+cM-2r\right)^{2}\right)}{4\left(M-r\right)^{4}}.$ (3.111)
Based on these results, we aim to understand the physical significance of the black hole solution (3.107) in DFT. In the framework of DFT, the entire doubled metric of spacetime is given by:
$ds^{2}=\underset{\mathrm{AdS\;vacuum\;in\;DFT}}{\underbrace{\underset{\mathrm{Black\;hole\;in\;ordinary\;string\;theory}}{\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left[-n\left(r\right)^{2}dt^{2}+B\left(r\right)dr^{2}+C\left(r\right)d\varphi^{2}\right]}}}+{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\left[B\left(r\right)^{-1}d\tilde{r}^{2}+C\left(r\right)^{-1}d\tilde{\varphi}^{2}\right]}}},$ (3.112)
This entire metric describes the AdS vacuum in double coordinates. The first bracket corresponds to the black hole solution known from ordinary string theory, while the second bracket represents the additional dimensions introduced by DFT. Since the first bracket is fully determined by the low-energy effective action, the usual interpretations in this framework can be directly applied. This suggests that this conventional, undoubled spacetime $x$ is self-contained and exhibits its own dynamics. If we consider the additional part $\left(\tilde{r},\tilde{\varphi}\right)$, there is no action to describe its geometry independently. Therefore, we must consider the entire doubled spacetime. DFT governs its dynamics, and it manifests as an AdS vacuum. In other words, if we obscure the doubled coordinates and focus only on the conventional region, the black hole and its singularity emerge as solutions of the low-energy effective action, as depicted in Figure. (1).
Doubled spacetime $\left(x,\tilde{x}\right)$ AdS vacuum Ordinary spacetime $(x)$ Black holes completely described by the low energy effective action
Figure 1: The dark blue region represents the conventional undoubled spacetime described by the low-energy effective action, which contains the black holes. When the additional region from the double coordinates is included, only DFT can describe the entire doubled spacetime, consistently characterizing it as an AdS vacuum.
To specify, we examine the following double black string solutions with specific values for $a$, $b$ and $c$:
1. When $a=1$, $b=0$, $c=1$, we obtain:
$ds^{2}=\left[-\left(1-\frac{M}{r}\right)dt^{2}+\frac{1}{r\left(r-M\right)}\frac{1}{\lambda^{2}}dr^{2}+\frac{1}{\lambda^{2}}d\varphi^{2}\right]+\left[r\left(r-M\right)\lambda^{2}d\tilde{r}^{2}+\lambda^{2}d\tilde{\varphi}^{2}\right]$ (3.113)
If we only consider the first square bracket, the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$, indicating this black hole solution possesses a spacetime singularity: $r=0$ where the event horizon is located at $r=M$. Considering the whole metric, there are two time-like directions $dr^{2}$ and $d\tilde{r}^{2}$ acrossing the event horizon. Additionally, accounting for the coordinates $\left(\tilde{r},\tilde{\varphi}\right)$, there is a coordinate singularity as $r\rightarrow\infty$. However, it is impossible to measure the curvature in the region $\left(\tilde{r},\tilde{\varphi}\right)$ independently. When considering the doubled coordinates, the metric corresponds to the AdS vacuum.
2. When $a=-1$, $b=0$, $c=1$, we have:
$ds^{2}=\left[-\left(1-\frac{M}{r}\right)^{-1}dt^{2}+\frac{1}{r\left(r-M\right)}\frac{1}{\lambda^{2}}dr^{2}+\frac{1}{\lambda^{2}}d\varphi^{2}\right]+\left[r\left(r-M\right)\lambda^{2}d\tilde{r}^{2}+\lambda^{2}d\tilde{\varphi}^{2}\right],\qquad M\leq 0,$ (3.114)
Considering the first square bracket, it is the solution of low-energy effective action and the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. This solution has a single curvature singularity at $r=M$, with a coordinate singularity as $r\rightarrow\infty$. And for complete metric, the metric still corresponds to the AdS vacuum.
3. When $a=0$, $b=\pm 1$, $c=1$, we have:
$ds^{2}$ $=$ $\left[-dt^{2}+\left(1-\frac{M}{r}\right)^{-1}\frac{1}{\lambda^{2}r^{2}}dr^{2}+\left(1-\frac{M}{r}\right)^{\pm 1}\frac{1}{\lambda^{2}}d\varphi^{2}\right]$ (3.115)
$+\left[\left(1-\frac{M}{r}\right)\lambda^{2}r^{2}d\tilde{r}^{2}+\left(1-\frac{M}{r}\right)^{\mp 1}\lambda^{2}d\tilde{\varphi}^{2}\right],\qquad M\geq 0,$
Here, $b=1$ for $M\geq 0$ and $b=-1$ for $M\leq 0$. When $b=-1$, there is a conical singularity at $r=M$ due to $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. When $b=1$, the curvature singularity is at $r=0$ because $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$. Furthermore, when the doubled coordinates are considered, the metric corresponds to the AdS vacuum.
4. When $a=b=c=0$, the solution is given as follows:
$ds^{2}$ $=$ $\left[-dt^{2}+\frac{1}{\lambda^{2}r^{2}}dr^{2}+\lambda^{\pm 2}d\varphi^{2}\right]+\left[\lambda^{2}r^{2}d\tilde{r}^{2}+\lambda^{\mp 2}d\tilde{\varphi}^{2}\right].$ (3.116)
The first square bracket of this metric describe flat spacetime, as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=0$. This implies that the half-geometry of the AdS vacuum in the doubled geometry can also manifest as flat spacetime when the doubled coordinates are ignored.
4 Conclusion and discussion
In this paper, we studied black hole solutions in DFT across two main sections. Initially, we directly calculated black hole solutions within standard DFT, where metrics involve two time-like directions. These solutions were categorized into three classes: two imposing to the strong constraint and one violating it by depending on dual coordinates. Due to the ambiguity in interpreting solutions with two time-like directions, further investigation is warranted for these cases. In the second part, we explored the space/time split DFT, where the theory only depends on conventional time-like directions, manifesting in a $1+2\left(D-1\right)$ dimensional spacetime $\left(t,x,\tilde{x}\right)$. In this framework, we investigated non-critical string theory incorporating a cosmological constant, also deriving solutions that violate the constraint. These solutions, similarly, are left for future investigation. Furthermore, we derived double black string solutions satisfying the strong constraint. In the context of the entire doubled spacetime, the generalized Ricci scalar consistently manifests as a negative constant, indicating AdS vacuum solutions. Upon hidden the dual coordinates $x$, the standard geometry described by $\tilde{x}$ reveals black hole solutions and curvature singularities.
Finally, we present some valuable study points for the near future:
Cosmology
Using our method, the cosmology of DFT can be revisited. By incorporating the cosmological constant, we can derive cosmological solutions. The corresponding generalized FLRW line element is given by:
$ds^{2}=-dt^{2}+a^{2}\left(t\right)dx^{i}dx_{i}+a^{-2}\left(t\right)d\tilde{x}^{i}d\tilde{x}_{i}.$ (4.117)
This solution addresses the double time problem discussed in reference [6]. Similar to the black hole solution, when considering the entire doubled spacetime, the generalized curvature implies that this solution corresponds to a dS vacuum.
Black hole soluions in Type II DFT
Considering the DFT action of type II theories [46, 47, 48]:
[…]
where the first term represents the standard DFT action describing the NS-NS sector, and the second term is the DFT action for the massless RR fields. The RR sector introduces two independent fields. The first field is $\mathbb{S}=\mathbb{S}^{\dagger}$, $\mathbb{S}\in\mathrm{Spin}\left(10,10\right)$. The second field is a spinor $\chi$, which can be represented by a Majorana-Weyl spinor of $O\left(10,10\right)$. Using our generating method, we can also easily obtain the double black $p$-brane solution of this action. It would be intriguing to explore whether this approach introduces new insights into the AdS/CFT correspondence.
Acknowledgements We are deeply indebt to Houwen Wu for many discussions and suggestions. This work was supported by NSFC Grant No.12105031, No.12347101 and PSFC Grant No. cstc2021jcyj-bshX0227.